### v1

# Beyond Maximum Likelihood: Variational Inequality Estimation for Generalized Linear Models

Abstract
Generalized linear models (GLMs) are fundamental tools for statistical modeling, with maximum likelihood estimation (MLE) serving as the classical method for parameter inference. While MLE performs well in canonical GLMs, it can become computationally inefficient near the true parameter value. In more general settings with non-canonical or fully general link functions, the resulting optimization landscape is often non-convex, non-smooth, and numerically unstable. To address these challenges, we investigate an alternative estimator based on solving the variational inequality (VI) formulation of the GLM likelihood equations, originally proposed by Juditsky and Nemirovski [juditsky2019signal] as an alternative for solving nonlinear least-squares problems. Unlike their focus on algorithmic convergence in monotone settings, we analyze the VI approach from a statistical perspective, comparing it systematically with the MLE. We also extend the theory of VI estimators to a broader class of link functions, including non-monotone cases satisfying a strong Minty condition, and show that it admits weaker smoothness requirements than MLE, enabling faster, more stable, and less locally trapped optimization. Theoretically, we establish both non-asymptotic estimation error bounds and asymptotic normality for the VI estimator, and further provide convergence guarantees for fixed-point and stochastic approximation algorithms. Numerical experiments show that the VI framework preserves the statistical efficiency of MLE while substantially extending its applicability to more challenging GLM settings.
1 Introduction
Generalized linear models (GLMs) are a cornerstone of statistics and machine learning, offering a unified framework that extends linear regression to accommodate diverse response types through the exponential-family formulation [nelder1972generalized, mccullagh1989generalized]. By linking the conditional mean of the response to a linear predictor via a specified link function, GLMs combine interpretability and flexibility, encompassing classical models such as linear, logistic, and Poisson regression as special cases. Their broad applicability has made GLMs indispensable across scientific, engineering, and data-driven decision-making domains.
The central inferential task in GLMs is to estimate the regression parameter characterizing the relationship between covariates and responses. The maximum likelihood estimator (MLE) has long served as the canonical approach, offering strong asymptotic guarantees such as efficiency and consistency under correct model specification. However, the MLE encounters limitations in modern modeling scenarios. When the link function departs from the canonical form or incorporates non-smooth or nonlinear structures designed for robustness and flexibility, the associated likelihood-based loss function can become highly ill-conditioned, non-convex, or non-differentiable. These issues often result in slow convergence, sensitivity to initialization, and numerical instability—challenges that are amplified in high-dimensional data settings. These difficulties motivate the exploration of alternative formulations that retain the interpretability of GLMs while improving computational efficiency and providing stronger theoretical guarantees.
In this paper, we study variational inequality (VI)-based estimation as a principled alternative to maximum likelihood estimation for GLMs. The VI framework, initially proposed by Juditsky and Nemirovski [juditsky2019signal] as an alternative for solving nonlinear least-squares problems, provides a flexible operator-based formulation of estimation. Unlike the MLE, which minimizes the negative log-likelihood, the VI estimator is defined through an equilibrium condition that generalizes first-order methods based on the MLE. This perspective naturally accommodates non-canonical and even non-monotone link functions, relaxing the smoothness and convexity requirements inherent in likelihood-based estimation. The proposed framework unifies likelihood-based estimation and variational inequalities under a common operator perspective, enabling rigorous analysis of both statistical accuracy and algorithmic convergence. Our work contributes to the growing literature on variational formulations of statistical estimation by establishing a new theoretical connection between GLMs and stochastic variational inequalities, along with corresponding statistical and computational guarantees.
Our analysis reveals that the VI formulation preserves key statistical properties of the MLE while offering stronger numerical stability and broader applicability. We establish both non-asymptotic estimation error bounds and asymptotic normality for the VI estimator, demonstrating that it achieves the optimal $\mathcal{O}(N^{-1/2})$ rate and a sandwich-type covariance structure analogous to that of the MLE. Furthermore, we provide convergence guarantees for fixed-point and stochastic approximation algorithms used to compute the estimator. Empirical results confirm that the VI estimator achieves comparable or superior accuracy and faster convergence than MLE in the finite sample regime, particularly with limited sample sizes, and for GLMs with non-canonical or non-smooth link functions.
The remainder of the paper is organized as follows. Section 2 introduces the problem setup and the proposed VI formulation that generalizes the MLE. Section 3 develops the theoretical relationship between VI and MLE under canonical and non-canonical link functions. Section 4 presents the main statistical results, including finite-sample error bounds and asymptotic normality. Section 5 analyzes the convergence behavior of fixed-point and stochastic approximation algorithms for solving the empirical VI problem. Section 6 reports numerical experiments comparing VI and MLE across various GLM settings, and Section 7 concludes with remarks on designing link functions for improved statistical and computational performance. Additional proofs and derivations are provided in the Appendix.
1.1 Notation
The notation in the paper is standard. Let $\mathbb{R}^{d}$ denote the $d$-dimensional Euclidean space equipped with the inner product $\langle\bm{u},\bm{v}\rangle:=\bm{u}^{\top}\bm{v}$ and the induced norm $\|\bm{u}\|:=\sqrt{\langle\bm{u},\bm{u}\rangle}$. We write $\|\cdot\|_{\infty}$ for the infinity norm. For a convex set $\mathcal{B}\subseteq\mathbb{R}^{d}$, the Euclidean projection of a point $\bm{\beta}$ onto $\mathcal{B}$ is defined as $\operatorname{proj}_{\mathcal{B}}(\bm{\beta}):=\operatorname*{argmin}_{\bm{\beta}^{\prime}\in\mathcal{B}}\|\bm{\beta}^{\prime}-\bm{\beta}\|$, and the corresponding distance is given by $\operatorname{dist}(\bm{\beta},\mathcal{B}):=\|\bm{\beta}-\operatorname{proj}_{\mathcal{B}}(\bm{\beta})\|$. For a general vector field $F:\mathbb{R}^{d}\to\mathbb{R}^{d}$, its solution set is defined as
$\operatorname{Sol}(F):=\left\{\hat{\bm{\beta}}\in\mathbb{R}^{d}:\langle F(\hat{\bm{\beta}}),\bm{\beta}-\hat{\bm{\beta}}\rangle\geq 0,\ \forall\bm{\beta}\in\mathbb{R}^{d}\right\}=\left\{\hat{\bm{\beta}}\in\mathbb{R}^{d}:F(\hat{\bm{\beta}})=\mathbf{0}\right\}.$
[…]
2 Background and Problem Setup
We begin by reviewing the GLM framework and establishing notation used throughout the paper. We then introduce the MLE approach and its VI reformulation, which serves as the foundation for our theoretical analysis.
Given i.i.d. samples $(\bm{x},y)$, where $\bm{x}\in\mathbb{R}^{d}$ denotes the vector of regressors and $y\in\mathbb{R}$ denotes the response, GLM specifies the conditional distribution of response $y\mid\bm{x}$ using the exponential family. The mean response is related to a linear predictor through a specified link function $g:\mathbb{R}\to\mathbb{R}$:
$\mathbb{E}[y\mid\bm{x}]=g^{-1}\left(\beta_{0}+\sum_{j=1}^{d}\beta_{j}x_{j}\right),$ (2.1)
where $\bm{\beta}=(\beta_{0},\dots,\beta_{d})\in\mathbb{R}^{d+1}$ is the coefficient vector with the intercept term. GLMs include many classical models as special cases. For instance, logistic regression corresponds to a GLM with a binomial distribution and a logit link, while Poisson regression corresponds to a GLM with a Poisson distribution and a logarithm link function.
A standard approach for estimating $\bm{\beta}$ is the MLE, which maximizes the log-likelihood of the observed data. Given a statistical model with probability density (or mass) function $p(y;\bm{\beta})$ and a dataset $\{y^{i}\}_{i=1}^{N}$, the MLE is defined as
$\max_{\bm{\beta}\in\mathbb{R}^{d+1}}\;\prod_{i=1}^{N}p(y^{i};\bm{\beta}),$
which is equivalently obtained by minimizing the empirical negative log-likelihood (NLL):
$\min_{\bm{\beta}\in\mathbb{R}^{d+1}}\;\sum_{i=1}^{N}-\log p(y^{i};\bm{\beta}).$
In the GLM framework, this objective can be expressed in the general empirical risk form with loss function $\ell$ and link function $g$:
$\min_{\bm{\beta}\in\mathbb{R}^{d+1}}\;\mathcal{L}_{N}(\bm{\beta}):=\sum_{i=1}^{N}\ell\left(g^{-1}\left(\beta_{0}+\sum_{j=1}^{d}\beta_{j}x^{i}_{j}\right),y^{i}\right).$ (MLE)
A variety of numerical algorithms can be employed to solve (MLE). Classical second-order methods, including Newton-type schemes such as Fisher scoring and iteratively reweighted least squares (IRLS) [green1984iteratively, mccullagh1989generalized], are commonly used in GLMs, while large-scale problems are often addressed using first-order approaches such as gradient descent or its stochastic variants [nesterov2013introductory, kushner2003stochastic].
In classical settings such as linear, logistic, and Poisson regression with canonical link functions, the optimization problem (MLE) is convex and smooth, ensuring global optimality and fast convergence of standard algorithms. However, when the GLM employs a non-canonical link, the optimization landscape can become ill-posed, leading to slow convergence of these numerical methods; see, e.g., [he2020point]. In addition, if the inverse link function is non-monotone or non-smooth (see, e.g., [cao2015poisson], where in practice one has to impose a lower bound on the Poisson intensity function), the resulting NLL objective is generally non-convex and non-smooth. Such situations arise in practice, for example, in Poisson regression with a clipped exponential link or in non-canonical extensions of GLMs that use piecewise linear activations such as hinge or ReLU functions. To address these challenges, it is natural to consider reformulating the estimation problem from an operator perspective, one that captures the equilibrium condition defining the MLE but relaxes convexity and differentiability requirements.
To overcome these difficulties, we reformulate the estimation problem as a variational inequality. The VI framework introduces an operator-based characterization that allows analysis and computation under weaker regularity assumptions. Let $V:\mathbb{R}^{d+1}\to\mathbb{R}^{d+1}$ denote the population operator
$V(\bm{\beta}):=\mathbb{E}_{(\bm{x},y)\sim\mathbb{P}}\!\left[\left(g^{-1}\left(\beta_{0}+\sum_{j=1}^{d}\beta_{j}x_{j}\right)-y\right)\tilde{\bm{x}}\right$],
where we introduce the augmented feature vector $\tilde{\bm{x}}:=[1;\bm{x}]\in\mathbb{R}^{d+1}$. The target parameter $\bm{\beta}^{\star}$ is defined as the solution to the variational inequality
$\langle V(\bm{\beta}^{\star}),\,\bm{\beta}-\bm{\beta}^{\star}\rangle\geq 0,\quad\forall\,\bm{\beta}\in\mathbb{R}^{d+1}.$ (2.2)
Given i.i.d. observations, $\{(\bm{x}^{i},y^{i})\}_{i=1}^{N}$, the empirical VI estimator $\hat{\bm{\beta}}_{N}$ is obtained by the empirical average:
$\hat{\bm{\beta}}_{N}\in\left\{\hat{\bm{\beta}}\in\mathbb{R}^{d+1}:\langle V_{N}(\hat{\bm{\beta}}),\bm{\beta}-\hat{\bm{\beta}}\rangle\geq 0,\ \forall\bm{\beta}\in\mathbb{R}^{d+1}\right\},$ (VI)
where
$V_{N}(\bm{\beta}):=\frac{1}{N}\sum_{i=1}^{N}\!\left(g^{-1}\left(\beta_{0}+\sum_{j=1}^{d}\beta_{j}x^{i}_{j}\right)-y^{i}\right)\tilde{\bm{x}}^{i}.$ (2.3)
The VI estimator generalizes the first-order optimality condition of the MLE and reduces to it under canonical link functions, implying that standard GLM estimation can be viewed as an implicit VI problem. This formulation was first considered by Juditsky and Nemirovski [juditsky2019signal] in the context of nonlinear least-squares estimation for GLMs. Their approach builds upon the seminal ideas of Rosenblatt’s perceptron algorithm [rosenblatt1958perceptron] and subsequent perceptron-like methods [kalai2009isotron, kakade2011efficient]. Since then, the VI framework has been extended to a variety of settings, demonstrating its versatility and effectiveness [juditsky2020convex, juditsky2023generalized, cheng2024point, zhou2024nonlinear, lou2025accurate]. Compared with the nonlinear least-squares formulation, which often leads to non-convex objectives under nonlinear link functions, the VI approach assumes only that the link function is strongly monotone, thereby inducing a strongly convex optimization structure and enabling stable computation. Building upon these foundational developments, our work provides a unified theoretical and algorithmic treatment of VI-based estimation for GLMs, establishing its statistical optimality and computational convergence properties.
2.1 Summary of main results
We develop a rigorous theoretical and algorithmic analysis of the proposed VI-based estimator for GLMs, establishing both statistical guarantees and computational convergence results. In contrast to existing results such as [juditsky2019signal], which primarily focus on using stochastic approximation schemes based on VI to solve nonlinear least-squares problems, our work highlights the statistical advantages and theoretical robustness of a VI-based estimator relative to the MLE for GLMs. The main contributions are summarized as follows:
  • Statistical guarantees. Beyond the non-asymptotic convergence results established for stochastic approximation schemes, we derive both finite-sample estimation error bounds and asymptotic normality for the sample-average approximation solution of the VI, extending the stochastic VI framework of [king1993asymptotic, xu2010sample]. Specifically, the estimator achieves the optimal finite-sample rate and asymptotic normality:
    $\|\hat{\bm{\beta}}_{N}-\bm{\beta}^{\star}\|\leq\tilde{\mathcal{O}}\left(\sqrt{\frac{d}{N}}\right)\ \mbox{w.h.p},\quad\text{and}\quad\sqrt{N}(\hat{\bm{\beta}}_{N}-\bm{\beta}^{\star})\ \xrightarrow{d}\ \mathcal{N}(\mathbf{0},\Sigma^{\star}),$
    for some covariance matrix $\Sigma^{\star}$ that depends on the true parameter and the data distribution. We also establish linear convergence of fixed-point iterative algorithms for solving the empirical VI and the standard sublinear convergence rate for stochastic approximation under streaming data.
  • Extension to general link functions. We extend the theoretical results within the stochastic VI framework to accommodate non-smooth, potentially non-monotone link functions that satisfy the strong Minty condition, such as ReLU and clipped links. These link functions can be used to handle, for instance, heavy-tailed data, stabilize variance, and improve robustness to outliers, thereby broadening the applicability of the proposed framework beyond the assumptions of existing methods. In contrast, achieving comparable performance with MLE requires the gradient of the composite mapping—i.e., the loss function composed with the inverse link—to satisfy a gradient-based strong Minty condition, which involves an additional application of the chain rule. For models with non-canonical link functions, this condition is often difficult to satisfy, resulting in a flatter optimization landscape and slower convergence. When non-smooth or non-convex structures are involved, MLE algorithms may further become trapped in suboptimal solutions and suffer from numerical instability due to the failure of the chain rule.
3 From MLE to VI
We consider the GLM defined in (2.1), where the response variable follows an exponential-family distribution:
$y\mid\bm{x}\sim\mathrm{Exponential\,Family}(\mu),\quad\mu=g^{-1}(z),\quad z=\tilde{\bm{x}}^{\top}\bm{\beta}.$
In the formulation (MLE), the specification of (i) the loss function $\ell$ and (ii) the inverse link function $g^{-1}$ arises from modeling assumptions about the data-generating process.
3.1 Equivalence under canonical links
When the model employs a canonical link function, the gradient of the NLL function in (MLE) coincides exactly with the operator defining the variational inequality in (VI), i.e.,
$\nabla\mathcal{L}_{N}(\bm{\beta})=V_{N}(\bm{\beta}).$ (3.1)
This equivalence has been noted in [juditsky2019signal] for logistic regression. For instance, consider binary logistic regression, where the empirical NLL is given by
$\mathcal{L}_{N}(\bm{\beta})=-\frac{1}{N}\sum_{i=1}^{N}\left(y^{i}\log g^{-1}(\bm{\beta}^{\top}\tilde{\bm{x}}^{i})+(1-y^{i})\log(1-g^{-1}(\bm{\beta}^{\top}\tilde{\bm{x}}^{i}))\right),$
with gradient given by
$\nabla\mathcal{L}_{N}(\bm{\beta})=\frac{1}{N}\sum_{i=1}^{N}\frac{(g^{-1})^{\prime}(\bm{\beta}^{\top}\tilde{\bm{x}}^{i})}{g^{-1}(\bm{\beta}^{\top}\tilde{\bm{x}}^{i})\,(1-g^{-1}(\bm{\beta}^{\top}\tilde{\bm{x}}^{i}))}\left(g^{-1}(\bm{\beta}^{\top}\tilde{\bm{x}}^{i})-y^{i}\right)\tilde{\bm{x}}^{i}.$
For the logistic (sigmoid) link function $g^{-1}(z)=1/(1+e^{-z}),$ one can show that $(g^{-1})^{\prime}(z)=g^{-1}(z)(1-g^{-1}(z))$, which simplifies the gradient to
$\nabla\mathcal{L}_{N}(\bm{\beta})=\frac{1}{N}\sum_{i=1}^{N}\left(g^{-1}(\bm{\beta}^{\top}\tilde{\bm{x}}^{i})-y^{i}\right)\tilde{\bm{x}}^{i},$
and the last term above is identical to $V_{N}(\bm{\beta})$ in this case.
Such an equivalence also holds for linear, exponential, and Poisson regression models, indicating that under canonical link functions, the MLE and VI formulations share identical first-order optimality conditions. In contrast, when a non-canonical link function is used, the gradient of the likelihood and the VI operator differ, leading to distinct estimation dynamics and potentially different convergence behaviors. The similar identity holds for other types of link functions such as linear, Poisson, and exponential; representative examples are summarized in Table 1 and the derivations can be found in Appendix A.
Table 1: MLE and VI Formulations under Different Links
Model Loss Function $\ell(u,y)$ Inverse Link $g^{-1}(z)$ $\nabla\mathcal{L}(\bm{\beta})=V(\bm{\beta})?$
Logistic $-y\log u-(1-y)\log(1-u)$ $e^{z}/(1+e^{z})$
$1/2+\arctan(z)/\pi$
Normal $\frac{1}{2}(u-y)^{2}$ $z$
$\max\{0,z\}$
Exponential $\log u+y/u$ $z^{-1}$
$e^{-z}$
Poisson $-y\log u+u$ $e^{z}$
$\log(1+e^{z})$
$\max\{c,\min\{e^{z},C\}\}$
In the following non-canonical settings, the MLE landscape can become very flat, leading to slow convergence. Moreover, the loss function may be non-smooth or non-convex, violating standard subdifferential calculus rules [li2020understanding] and introducing additional challenges for both optimization and the statistical consistency of the MLE. Such cases arise, for example, in Poisson regression models with softplus or clipped exponential link functions.
3.2 Flatness of MLE optimization landscape
It is well known that MLE optimization algorithms can converge slowly near the true parameter, especially when the likelihood surface is nearly flat or ill-conditioned. We show that even when standard smooth link functions are applied to improve curvature, the resulting optimization landscape of the MLE can still exhibit considerably slower convergence than the VI formulation. This phenomenon is illustrated using the softplus link function, under which the MLE score and the VI operator define different estimating equations.
Recall the softplus inverse link function:
$g^{-1}(z)=\log(1+e^{z}),\quad(g^{-1})^{\prime}(z)=\sigma(z):=\frac{1}{1+e^{-z}},\quad(g^{-1})^{\prime\prime}(z)=\sigma(z)\left(1-\sigma(z)\right).$
For a single sample $(\bm{x},y)$, the Poisson negative log-likelihood is
$\mathcal{L}_{(\bm{x},y)}(\bm{\beta})=-y\log(g^{-1}(z))+g^{-1}(z),\quad z=\tilde{\bm{x}}^{\top}\bm{\beta},$
with gradient and Hessian given by, respectively,
$\nabla\mathcal{L}_{(\bm{x},y)}(\bm{\beta})=\sigma(z)\!\left(1-\frac{y}{\log(1+e^{z})}\right)\tilde{\bm{x}}$,
$\nabla^{2}\mathcal{L}_{(\bm{x},y)}(\bm{\beta})=\Bigg\{\sigma(z)\bigl(1-\sigma(z)\bigr)\!\left(1-\frac{y}{\log(1+e^{z})}\right)+\frac{y\,\sigma(z)^{2}}{\log(1+e^{z})^{2}}\Bigg\}\,\tilde{\bm{x}}\tilde{\bm{x}}^{\top}$.
On the other hand, the VI estimating operator is $V_{(\bm{x},y)}(\bm{\beta})=\left(g^{-1}(z)-y\right)\tilde{\bm{x}}$. At the population level, it follows that
$\nabla V(\bm{\beta})=\mathbb{E}_{\bm{x}\sim\mathbb{P}_{\bm{x}}}\left[\sigma(z)\cdot\tilde{\bm{x}}\tilde{\bm{x}}^{\top}\right],\quad\nabla^{2}\mathcal{L}(\bm{\beta})=\mathbb{E}_{\bm{x}\sim\mathbb{P}_{\bm{x}}}\left[\frac{\sigma(z)^{2}}{g^{-1}(z)}\cdot\tilde{\bm{x}}\tilde{\bm{x}}^{\top}\right$].
Since $g^{-1}(z)\geq\sigma(z)$ for all $z$ (with equality only as $z\to$-$\infty$), we have $\sigma(z)\geq\sigma(z)^{2}/g^{-1}(z)$ pointwise and hence
$\nabla V(\bm{\beta})-\nabla^{2}\mathcal{L}(\bm{\beta})=\mathbb{E}_{\bm{x}\sim\mathbb{P}_{\bm{x}}}\left[\left(\sigma(z)-\frac{\sigma(z)^{2}}{g^{-1}(z)}\right)\cdot\tilde{\bm{x}}\tilde{\bm{x}}^{\top}\right]\succeq\mathbf{0}.$
Moreover, we know that
$\frac{\sigma(z)}{\sigma^{2}(z)/g^{-1}(z)}=\frac{g^{-1}(z)}{\sigma(z)}=\frac{\log(1+e^{z})(1+e^{z})}{e^{z}}\rightarrow+\infty\quad\text{as}\quad z\to+\infty,$
which reveals a flat-growth regime in the right tail when the features are not linearly dependent (i.e., $\mathbb{E}_{\bm{x}\sim\mathbb{P}_{\bm{x}}}[\tilde{\bm{x}}\tilde{\bm{x}}^{\top}]\succ\mathbf{0}$): the curvature weight of the MLE decays relative to that of the VI formulation. Consequently, optimization solvers exhibit weaker local contraction when applied to the MLE, and statistical concentration bounds that depend on the curvature modulus become looser.
3.3 Dilemma of MLE with general links: Nonconvexity and nonsmoothness
Next, we show that the optimization problem associated with the MLE can become challenging in the presence of non-convex or non-smooth structures. We illustrate the difficulty using the two-sided clipped exponential link
$g^{-1}(z)=\max\{c,\,\min\{e^{z},\,C\}\},\qquad 0<c<C,\qquad z_{c}:=\log c,\ \ z_{C}:=\log C,$
and the Poisson per-sample negative log-likelihood
$(\ell\circ g^{-1})(z)=-\,y\log(g^{-1}(z))+g^{-1}(z).$
Then its piecewise form and derivatives are
$(\ell\circ g^{-1})(z)=\begin{cases}-\,y\log c+c,&z\leq z_{c},\\[2.0pt] -\,yz+e^{z},&z_{c}<z<z_{C},\\[2.0pt] -\,y\log C+C,&z\geq z_{C},\end{cases}\qquad(\ell\circ g^{-1})^{\prime}(z)=\begin{cases}0,&z<z_{c},\\[2.0pt] -\,y+e^{z},&z_{c}<z<z_{C},\\[2.0pt] 0,&z>z_{C}.\end{cases}$
Hence, the one-sided derivatives at the two kinks are
$(\ell\circ g^{-1})^{\prime}_{-}(z_{c})=0,\quad(\ell\circ g^{-1})^{\prime}_{+}(z_{c})=-\,y+c;\qquad(\ell\circ g^{-1})^{\prime}_{-}(z_{C})=-\,y+C,\quad(\ell\circ g^{-1})^{\prime}_{+}(z_{C})=0.$
Here $(\ell\circ g^{-1})^{\prime}_{-}$ and $(\ell\circ g^{-1})^{\prime}_{+}$ denote the left and right derivatives. Since a continuous piecewise-$C^{1}$ function on $\mathbb{R}$ being convex requires
[…]
and
$(\ell\circ g^{-1})^{\prime}_{-}(z_{C})\leq(\ell\circ g^{-1})^{\prime}_{+}(z_{C})\ \Longleftrightarrow\ -\,y+C\leq 0\ \Longleftrightarrow\ y\geq C,$
these conditions cannot hold simultaneously when $0<c<C$. Therefore, with two-sided clipping, the composition $\ell\circ g^{-1}$ is globally non-convex and non-smooth, with nondifferentiable points at $z=z_{c}$ and $z=z_{C}$. In particular, letting $y=1$ and $\bm{x}=\mathbf{0}$, we obtain
$\mathcal{L}(\beta)=(\ell\circ g^{-1})(\beta)=-\log g^{-1}(\beta)+g^{-1}(\beta).$
((a)) Function Value
((b)) Derivative
Figure 1: MLE loss function $\mathcal{L}(\beta)=-\log g^{-1}(\beta)+g^{-1}(\beta)$ and its derivative.
From Figure 1, we observe that the MLE objective is smooth only in the blue case ($c=0$, $C=1$). In the other two settings, the objective becomes non-convex and non-smooth: the gradient-based method can easily get stuck on the left and right plateaus in the orange case ($c=0.5$, $C=3$) and on the right plateau in the green case ($c=1$, $C=5$), failing to reach the global optimum at $\beta=0$. In contrast, Figure 2 shows that the corresponding VI vector field remains monotone in all three cases, suggesting that standard iterative methods can still converge to the global solution.
Figure 2: VI vector field $V(\beta)=g^{-1}(\beta)-1$.
4 Statistical Guarantees of VI Estimators
Having established the formulation in (VI) and its motivation, we now turn to the theoretical analysis of the proposed estimator’s statistical properties. This section derives both finite-sample and asymptotic guarantees for the empirical VI solution, defined as
$\hat{\bm{\beta}}_{N}\in\operatorname{Sol}(V_{N})=\left\{\bm{\beta}\in\mathbb{R}^{d+1}:\,V_{N}(\bm{\beta})=\sum_{i=1}^{N}\big(g^{-1}(\tilde{\bm{x}}^{i\top}\bm{\beta})-y^{i}\big)\tilde{\bm{x}}^{i}=\mathbf{0}\right\}.$
We begin with the following assumptions, which are used throughout the paper:
Assumption 4.1.
  1. (i)
    The inverse link function $g^{-1}$ is $L$-Lipschitz continuous.
  2. (ii)
    The true parameter $\bm{\beta}^{\star}$ and all the data $(\bm{x},y)$ satisfy
    $\left|g^{-1}\left(\beta^{\star}_{0}+\sum_{j=1}^{d}\beta^{\star}_{j}x_{j}\right)-y\right|\leq R.$
  3. (iii)
    The feature vector $\bm{x}$ satisfies $|x_{j}|\leq M$ and $|x_{j}\cdot y|\leq M$ for all $j\in[d]$ and all $(\bm{x},y)$.
  4. (iv)
    The population operator $V$ and the empirical operator $V_{N}$ satisfy the strong Minty condition with modulus $\mu>0$:
    $\langle V(\bm{\beta}),\bm{\beta}-\bm{\beta}^{\star}\rangle\geq\mu\|\bm{\beta}-\bm{\beta}^{\star}\|^{2},\quad\langle V_{N}(\bm{\beta}),\bm{\beta}-\hat{\bm{\beta}}_{N}\rangle\geq\mu\|\bm{\beta}-\hat{\bm{\beta}}_{N}\|^{2},\quad\forall\bm{\beta}\in\mathbb{R}^{d+1}.$
    Consequently, the solution sets $\operatorname{Sol}(V)$ and $\operatorname{Sol}(V_{N})$ are nonempty and singleton.
In the above assumption, the only nontrivial requirement is the strong Minty condition. The following lemma shows that the vector field $V_{N}$ satisfies the strong Minty condition when the inverse link function $g^{-1}$ satisfies an averaged Minty’s condition (which can be implied by the strong monotonicity). With a fixed sample size $N$, we collect the covariates into the data matrix $\bm{X}_{N}\in\mathbb{R}^{N\times d}$ defined as
$\bm{X}_{N}:=\begin{bmatrix}x^{1}_{1}&\cdots&x^{1}_{d}\\ \vdots&\ddots&\vdots\\ x^{N}_{1}&\cdots&x^{N}_{d}\end{bmatrix}$,
where $x^{i}_{j}$ denotes the $j$-th covariate of the $i$-th observation. A population version can be derived under analogous conditions.
Lemma 4.1 (Sufficient Condition for Strong Minty Condition).
Suppose that the inverse link function $g^{-1}$ satisfies the strong monotonicity with modulus $\mu_{g}\geq 0$ or the averaged strong Minty condition with $\operatorname{Sol}(V_{N})\neq\emptyset$ that
$\frac{1}{N}\sum_{i=1}^{N}\left(g^{-1}(\tilde{\bm{x}}^{i\top}\bm{\beta})-y^{i}\right)\cdot\tilde{\bm{x}}^{i\top}(\bm{\beta}-\operatorname{proj}_{\operatorname{Sol}(V_{N})}(\bm{\beta}))\geq\frac{\mu_{g}}{N}\cdot\sum_{i=1}^{N}\left|\tilde{\bm{x}}^{i\top}(\bm{\beta}-\operatorname{proj}_{\operatorname{Sol}(V_{N})}(\bm{\beta}))\right|^{2}.$ (4.1)
Then the vector field $V_{N}$ satisfies the strong Minty condition with modulus $\mu_{g}\sigma_{N}^{2}/N$, where $\sigma_{N}$ is the minimal singular value of $\tilde{\bm{X}}_{N}:=[\bm{1}\;\bm{X}_{N}]$, i.e.,
$\langle V_{N}(\bm{\beta}),\bm{\beta}-\operatorname{proj}_{\operatorname{Sol}(V_{N})}(\bm{\beta})\rangle\geq\frac{\mu_{g}\sigma_{N}^{2}}{N}\cdot\|\bm{\beta}-\operatorname{proj}_{\operatorname{Sol}(V_{N})}(\bm{\beta})\|^{2},\quad\forall\bm{\beta}\in\mathbb{R}^{d+1}.$
Remark 4.1 (Strong Minty condition).
The strong Minty condition is a relatively mild regularity requirement compared to strong monotonicity. In Assumption 4.1, we adopt a slightly stronger version to ensure the uniqueness of both the true parameter $\bm{\beta}^{\star}$ and the sample-averaged solution $\hat{\bm{\beta}}_{N}$ for simplicity. Nevertheless, this assumption can be relaxed to the standard projection-based form in Definition B.1, by replacing $\bm{\beta}^{\star}$ and $\hat{\bm{\beta}}_{N}$ with $\operatorname{proj}_{\operatorname{Sol}(V)}(\bm{\beta})$ and $\operatorname{proj}_{\operatorname{Sol}(V_{N})}(\bm{\beta})$, respectively, as also discussed in Lemma 4.1. Intuitively, it requires that the vector field $F$ exhibits at least linear growth away from the solution, with a uniform positive coefficient that prevents flat or degenerate regions around the solution. Geometrically, in the one-dimensional case, this means that $F$ defines a supporting hyperplane at the solution that strictly separates the solution from all other points, thereby ensuring the local stability of the solution. Further discussion and detailed analysis are provided in Appendix B; see also related treatments in [huang2023beyond].
With these preparations in place, we now proceed to derive finite-sample estimation error bounds that characterize the deviation between the empirical solution $\hat{\bm{\beta}}_{N}$ and the population parameter $\bm{\beta}^{\star}$ in Section 4.1. We then establish that the VI estimator is $\mathcal{O}(N^{-1/2})$-consistent and asymptotically normal, exhibiting a sandwich-type covariance structure analogous to that of the MLE, as detailed in Section 4.2.
4.1 Finite-sample estimation error
As a first step, we establish the following auxiliary lemma. Associated with the pair $(\bm{x},y)$, we recall the vector field $V_{(\bm{x},y)}:\mathbb{R}^{d+1}\rightarrow\mathbb{R}^{d+1}$
$V_{(\bm{x},y)}(\bm{\beta}):=\left(g^{-1}\left(\beta_{0}+\sum_{j=1}^{d}\beta_{j}x_{j}\right)-y\right)\tilde{\bm{x}}$. (4.2)
Lemma 4.2.
Suppose that Assumptions 4.1 (i)-(iii) hold. Then for every $(\bm{x},y)$ we have
$\|V_{(\bm{x},y)}(\bm{\beta})\|_{\infty}\leq RM,\quad\|V_{(\bm{x},y)}(\bm{\beta})\|\leq\sqrt{d+1}RM,$
and
$\|V_{(\bm{x},y)}(\bm{\beta}_{2})-V_{(\bm{x},y)}(\bm{\beta}_{1})\|\leq(LdM^{2}+L)\cdot\|\bm{\beta}_{2}-\bm{\beta}_{1}\|.$
Now, we are ready to derive the following estimation error of the variational inequality (VI).
Theorem 4.3 (Estimation error).
Suppose that Assumption 4.1 holds. Then for any $\epsilon\in(0,1)$, with the probability at least $1-\epsilon$, we have the following estimation error
$\|\hat{\bm{\beta}}_{N}-\bm{\beta}^{\star}\|\leq\frac{\|V_{N}(\bm{\beta}^{\star})\|}{\mu}\leq\frac{RM}{\mu}\sqrt{\frac{2(d+1)\ln(2(d+1)/\epsilon)}{N}}.$
Remark 4.2.
For the MLE, a similar finite-sample error bound can be derived as $\tilde{\mathcal{O}}(\sqrt{d}/(\mu^{\prime}\sqrt{N})),$ where $\mu^{\prime}>0$ denotes the curvature modulus of the composite mapping $\ell\circ g^{-1}$ that satisfies the restricted secant inequality (RSI), i.e.,
$\langle\nabla\mathcal{L}_{N}(\bm{\beta}),\bm{\beta}-\bm{\beta}^{\star}\rangle\geq\mu^{\prime}\|\bm{\beta}-\bm{\beta}^{\star}\|^{2},$
a condition stronger than the Polyak–Łojasiewicz inequality and weaker than the strong convexity; see [karimi2016linear]. Recall that the gradient of the empirical loss of (MLE) is given by
$\nabla\mathcal{L}_{N}(\bm{\beta})=\mathbb{E}_{(\bm{x},y)\sim\mathbb{P}_{N}}\!\left[(g^{-1})^{\prime}\!\left(\beta_{0}+\sum_{j=1}^{d}\beta_{j}x_{j}\right)\cdot\ell^{\prime}\!\left(g^{-1}\!\left(\beta_{0}+\sum_{j=1}^{d}\beta_{j}x_{j}\right),\,y\right)\cdot\tilde{\bm{x}}\,\right$].
If the RSI holds with constant $\mu^{\prime}$, then the maximum likelihood estimator $\bar{\bm{\beta}}_{N}$
$\mu^{\prime}\sigma^{2}\|\bar{\bm{\beta}}_{N}-\bm{\beta}^{\star}\|^{2}\leq\langle\nabla\mathcal{L}_{N}(\bm{\beta}^{\star}),\bm{\beta}^{\star}-\bar{\bm{\beta}}_{N}\rangle,$
which implies the bound
$\|\bar{\bm{\beta}}_{N}-\bm{\beta}^{\star}\|\leq\frac{\|\nabla\mathcal{L}_{N}(\bm{\beta}^{\star})\|}{\mu^{\prime}}.$
Compared with the proof of the VI estimator in Theorem 4.3, where $g^{-1}$ itself satisfies the strong Minty condition, the MLE requires the composite operator $(\ell^{\prime}\circ g^{-1})\cdot(g^{-1})^{\prime}$ to satisfy an analogous Minty-type condition. This composite structure makes the regularity requirement for MLE substantially more intricate.
4.2 Asymptotic normality of VI estimators
We now proceed to establish the asymptotic normality of the VI estimator. Our goal is to show that $\sqrt{N}(\hat{\bm{\beta}}_{N}-\bm{\beta}^{\star})$ converges in distribution to a normal random vector with mean zero and a sandwich-type covariance matrix. To handle potential nonsmoothness of the link function $g^{-1}$, we employ tools from empirical process theory and a local linearization argument around the population solution $\bm{\beta}^{\star}$.
Since $g^{-1}$ is globally Lipschitz by Assumption 4.1, Rademacher’s theorem [rockafellar2009variational, Theorem 9.60] implies that it is differentiable almost everywhere on $\mathbb{R}$, and hence its nondifferentiability set
$\mathcal{K}:=\{t:(g^{-1})^{\prime}(t)\text{ does not exist}\}$
has Lebesgue measure zero. We further assume the following condition.
Assumption 4.2.
The random variable $\langle\tilde{\bm{x}},\bm{\beta}^{\star}\rangle$ does not fall on the nondifferentiability set $\mathcal{K}$ of $g^{-1}$, that is,
$\mathbb{P}\big(\langle\tilde{\bm{x}},\bm{\beta}^{\star}\rangle\in\mathcal{K}\big)=0.$
This assumption is not restrictive in practice, since $\mathcal{K}$ has Lebesgue measure zero and thus the event $\langle\tilde{\bm{x}},\bm{\beta}^{\star}\rangle\in\mathcal{K}$ occurs with probability zero for any continuous feature distribution. With this mild condition, the population VI operator $V$ is guaranteed to be differentiable at the true parameter $\bm{\beta}^{\star}$.
Lemma 4.4.
Suppose that Assumptions 4.1 and 4.2 hold. Then the function $V$ is Fréchet differentiable at $\bm{\beta}^{\star}$ with its Jacobian
$\nabla V(\bm{\beta}^{\star})=\mathbb{E}_{(\bm{x},y)\sim\mathbb{P}}\left[(g^{-1})^{\prime}(\langle\tilde{\bm{x}},\bm{\beta}^{\star}\rangle)\,\tilde{\bm{x}}\tilde{\bm{x}}^{\top}\right]\in\mathbb{R}^{(d+1)\times(d+1)}.$
Owing to the averaging-induced smoothing effect, we obtain the asymptotic normality result.
Theorem 4.5 (Asymptotic normality of the VI estimator).
Suppose Assumptions 4.1 and 4.2 hold, and in addition
  1. (A1)
    The Jacobian $\nabla V(\bm{\beta}^{\star})$ is non-singular.
  2. (A2)
    The covariance matrix
    $\Gamma:=\mathbb{E}_{(\bm{x},y)\sim\mathbb{P}}[V_{(\bm{x},y)}(\bm{\beta}^{\star})V_{(\bm{x},y)}(\bm{\beta}^{\star})^{\top}]=\mathbb{E}_{(\bm{x},y)\sim\mathbb{P}}[(y-g^{-1}(\tilde{\bm{x}}^{\top}\bm{\beta}^{\star}))^{2}\tilde{\bm{x}}\tilde{\bm{x}}^{\top}]\in\mathbb{R}^{(d+1)\times(d+1)},$
    is finite and positive definite.
Then $\hat{\bm{\beta}}_{N}$ is $\sqrt{N}$-consistent and
$\sqrt{N}(\hat{\bm{\beta}}_{N}-\bm{\beta}^{\star})\ \xrightarrow{d}\ \mathcal{N}\!\big(0,\;\nabla V(\bm{\beta}^{\star})^{-1}\Gamma\nabla V(\bm{\beta}^{\star})^{-\top}\big).$
A classical result in asymptotic statistics states that the MLE is asymptotically efficient, namely, it is consistent, asymptotically unbiased, and attains the Cramér–Rao lower bound (CRB) under correct model specification. The asymptotic covariance of the proposed VI estimator,
$\Sigma_{\mathrm{VI}}=\nabla V(\bm{\beta}^{\star})^{-1}\Gamma\nabla V(\bm{\beta}^{\star})^{-\top},$
shares the same “sandwich” structure as that of the MLE,
$\Sigma_{\mathrm{MLE}}=\big(\nabla^{2}\mathcal{L}(\bm{\beta}^{\star})\big)^{-1}\operatorname{Cov}(\nabla\mathcal{L}_{(\bm{x},y)}(\bm{\beta}^{\star}))\big(\nabla^{2}\mathcal{L}(\bm{\beta}^{\star})\big)^{-\top},$
where the expected Hessian and the score covariance are given by
$\nabla^{2}\mathcal{L}(\bm{\beta}^{\star})=\mathbb{E}_{(\bm{x},y)\sim\mathbb{P}}\!\left[\frac{(g^{-1})^{\prime}(\tilde{\bm{x}}^{\top}\bm{\beta}^{\star})^{2}}{\operatorname{Var}(y\mid\tilde{\bm{x}})}\,\tilde{\bm{x}}\tilde{\bm{x}}^{\top}\right$],
and
$\operatorname{Cov}(\nabla\mathcal{L}_{(\bm{x},y)}(\bm{\beta}^{\star})):=\$$\mathbb{E}_{(\bm{x},y)\sim\mathbb{P}}\!\left[\nabla\mathcal{L}_{(\bm{x},y)}(\bm{\beta}^{\star})\nabla\mathcal{L}_{(\bm{x},y)}(\bm{\beta}^{\star})^{\top}\right]$
[…]
For correctly specified models, the information identity
$\operatorname{Cov}\!\big(\nabla\mathcal{L}_{(\bm{x},y)}(\bm{\beta}^{\star})\big)=\nabla^{2}\mathcal{L}(\bm{\beta}^{\star})=I(\bm{\beta}^{\star})$
holds, and the MLE achieves the Fisher information bound $I(\bm{\beta}^{\star})^{-1}$, attaining the CRB. In the following remark, we show that the MLE can indeed be asymptotically more efficient than the VI estimator under correct model specification. Their asymptotic covariances coincide only when the link function is canonical. This observation does not contradict our earlier discussion, as the VI estimator may still outperform the MLE in the presence of model or variance misspecification.
Remark 4.3 (Comparison of VI with MLE on asymptotic statistical efficiency).
For correctly specified models with general link functions, the VI estimator generally does not satisfy the corresponding identity $\Gamma=\nabla V(\bm{\beta}^{\star})$, since the residual covariance and the Jacobian $\nabla V(\bm{\beta}^{\star})$ coincide only for special canonical links. Let
$\Sigma_{\mathrm{MLE}}^{-1}=\mathbb{E}\!\left[\frac{(g^{-1})^{\prime}(\tilde{\bm{x}}^{\top}\bm{\beta}^{\star})^{2}}{\operatorname{Var}(y\mid\tilde{\bm{x}})}\,\tilde{\bm{x}}\tilde{\bm{x}}^{\top}\right],\quad\bm{r}_{1}=\sqrt{\operatorname{Var}(y\mid\tilde{\bm{x}})}\,\tilde{\bm{x}},\quad\bm{r}_{2}=\frac{(g^{-1})^{\prime}(\tilde{\bm{x}}^{\top}\bm{\beta}^{\star})}{\sqrt{\operatorname{Var}(y\mid\tilde{\bm{x}})}}\,\tilde{\bm{x}}$.
Then $\mathbb{E}[\bm{r}_{1}\bm{r}_{1}^{\top}]=\Gamma$, $\mathbb{E}[\bm{r}_{2}\bm{r}_{2}^{\top}]=\Sigma_{\mathrm{MLE}}^{-1}$, $\mathbb{E}[\bm{r}_{1}\bm{r}_{2}^{\top}]=\nabla V(\bm{\beta}^{\star})$, and hence the block moment matrix
$\begin{bmatrix}\Gamma&\nabla V(\bm{\beta}^{\star})\\[2.0pt] \nabla V(\bm{\beta}^{\star})&\Sigma_{\mathrm{MLE}}^{-1}\end{bmatrix}\succeq\mathbf{0}.$
By the Schur complement, this implies the MLE is never less efficient than the VI estimator, i.e.,
$\Sigma_{\mathrm{MLE}}\ \preceq\ \Sigma_{\mathrm{VI}}.$
The equality holds if and only if $\bm{r}_{1}$ and $\bm{r}_{2}$ are a.s. linearly dependent, i.e., up to a constant $c>0$,
$\mathbb{E}\big[(y-g^{-1}(\tilde{\bm{x}}^{\top}\bm{\beta}^{\star}))^{2}\mid\tilde{\bm{x}}\big]=c\cdot(g^{-1})^{\prime}(\tilde{\bm{x}}^{\top}\bm{\beta}^{\star})\quad\text{a.s. in }\tilde{\bm{x}}$,
which corresponds to the case where the link $g$ is canonical (up to a constant scale). Hence, under correct specification, the MLE dominates the VI estimator in efficiency, with equality attained exactly for canonical links. As a result, the VI estimator may exhibit a loss of statistical efficiency relative to the MLE, for which the MLE is asymptotically optimal in the sense of attaining the CRB. Nevertheless, when the model or variance component is misspecified, the VI formulation, which relies solely on the mean relation, remains consistent and may even achieve a smaller asymptotic variance for the mean component, providing a robust alternative to likelihood-based estimation.
5 Convergence Guarantees of VI-Based Algorithms
This section investigates the computational efficiency of solving the empirical VI problem using iterative algorithms. We begin by analyzing the convergence behavior of a fixed-point iterative method under a fixed sample size $N$. We then extend the analysis to the stochastic approximation setting, where the estimation is performed using a streaming data scheme.
For the following deterministic fixed-point iterative scheme under a fixed sample size, we establish its linear convergence rates under the strong Minty condition.
Input : Initialization $\bm{\beta}^{0}$, step size $\eta>0$, data $(\bm{x},y)\sim\mathbb{P}_{N}$
1 for $t=0,1,2,\ldots,T-1$ do
2 $\bm{\beta}^{t+1}:=\bm{\beta}^{t}-\eta\cdot V_{N}(\bm{\beta}^{t})$
3 end for
Algorithm 1 Fixed-Point Iterative Method
Theorem 5.1 (Linear convergence of fixed-point method).
Suppose that Assumption 4.1 holds with the step size
$\eta=\frac{\mu}{L^{2}(1+dM^{2})^{2}},$
where $\mu>0$ is the strong Minty modulus in Assumption 4.1. Then the sequence of estimates $\bm{\beta}^{t}$ given by the Algorithm 1 for every $t$ satisfies
$\|\bm{\beta}^{t}-\hat{\bm{\beta}}_{N}\|^{2}\leq\left(1-\frac{\mu^{2}}{L^{2}(1+dM^{2})^{2}}\right)^{t}\cdot\|\bm{\beta}^{0}-\hat{\bm{\beta}}_{N}\|^{2},\quad t=0,1,2,\ldots$
Furthermore, for any $\epsilon\in(0,1)$, with the probability at least $1-\epsilon$, we have the following estimation error
$\|\bm{\beta}^{t}-\bm{\beta}^{\star}\|\leq\left(1-\frac{\mu^{2}}{L^{2}(1+dM^{2})^{2}}\right)^{t/2}\cdot\|\bm{\beta}^{0}-\hat{\bm{\beta}}_{N}\|+\frac{RM}{\mu}\sqrt{\frac{2(d+1)\ln(2(d+1)/\epsilon)}{N}}.$
Next, we consider the stochastic approximation setting, in which the estimator is updated from streaming data.
Input : Initialization $\bm{\beta}^{0}$, step size $\eta^{t}>0$, streaming data $(\bm{x}^{t},y^{t})\sim\mathbb{P}$
1 for $t=0,1,2,\ldots,T-1$ do
2 $\bm{\beta}^{t+1}:=\bm{\beta}^{t}-\eta^{t}\cdot V_{(\bm{x}^{t},y^{t})}(\bm{\beta}^{t})$
3 end for
Algorithm 2 Stochastic Approximation
Theorem 5.2 (Sublinear estimation error of stochastic approximation).
Suppose that Assumption 4.1 holds. Then the sequence $\{\bm{\beta}^{t}\}_{t\geq 0}$ generated by Algorithm 2 with step size
$\eta^{t}=\frac{1}{\mu(t+1)},\quad t=0,1,2,\ldots,$
where $\mu>0$ is the strong Minty modulus in Assumption 4.1, satisfies
$\mathbb{E}\|\bm{\beta}^{t}-\bm{\beta}^{\star}\|^{2}\leq\frac{c_{0}(d+1)R^{2}M^{2}}{\mu^{2}(t+1)},\quad t=0,1,2,\ldots,$
for some constant $c_{0}>1$ that depends on the initialization.
Remark 5.1.
Theorem 5.2 parallels the stochastic approximation result of [juditsky2019signal], which established an $\mathcal{O}(1/t)$ mean-square convergence rate for the stochastic mirror-prox algorithm under Lipschitz continuity and strong monotonicity of the population operator. Our analysis adopts an analogous Lipschitz condition but replaces strong monotonicity with the weaker strong Minty condition, which accommodates potentially non-monotone operators induced by non-canonical link functions such as clipped or piecewise-linear activations in GLMs. Consequently, Theorem 5.2 extends classical stochastic gradient method convergence guarantees beyond monotone regimes to more general, possibly non-convex or non-smooth GLM settings, while preserving $\mathcal{O}(1/t)$ expected mean-square error rate.
These results collectively show that the proposed VI-based algorithms achieve both statistical consistency and efficient convergence, even under non-smooth or non-monotone link functions, thereby underscoring the computational advantages of the proposed framework for general estimation problems.
6 Numerical Experiments
We examine the finite-sample performance of the VI estimator relative to the MLE in Poisson regression with various link functions. For each experiment, the true parameter vector is fixed as $\bm{\beta}^{\star}=d^{-1/2}(1,\dots,1)\in\mathbb{R}^{d}$, and the regressors $\bm{x}^{i}\in\mathbb{R}^{d}$, $i=1,\dots,N$, are drawn i.i.d. from $\mathcal{N}(0,I_{d})$. Given $\bm{x}^{i}$, the response $y^{i}\in\mathbb{R}$ is linked to the linear predictor $\bm{\beta}^{\top}\bm{x}^{i}$ without an intercept term, and is generated according to
$y^{i}\sim\mathrm{Poisson}\left(g^{-1}(\bm{\beta}^{\top}\bm{x}^{i})\right),$
where the inverse link function $g^{-1}$ is chosen from the following:
  • log: $g^{-1}(z)=e^{z}$,
  • softplus: $g^{-1}(z)=\log(1+e^{z})$,
  • clipped exponential: $g^{-1}(z)=\min\{e^{z},2\}$,
  • scaled Gaussian-mixture CDF with two components: $g^{-1}(z)=1.65\cdot\Phi\big(\frac{z+0.5}{0.7}\big)+1.35\cdot\Phi\big(\frac{z-1.2}{0.5}\big)$,
where $\Phi(z)=\int_{-\infty}^{z}\exp(-x^{2}/2)/\sqrt{2\pi}dx$ is the standard Gaussian CDF. The log link corresponds to the canonical Poisson regression model, in which the VI and MLE updates coincide. The softplus and clipped exponential links, respectively, introduce bounded curvature and truncation in the mean function. We also considered the Gaussian-mixture CDF as a link function, which provides a non-convex example intended to test the robustness in more challenging settings.
For each link, both estimators are computed by iterative updates of the form
$\bm{\beta}^{k+1}$ $=\bm{\beta}^{k}-\eta^{k}V(\bm{\beta}^{k}),$
$\bm{\beta}^{k+1}$ $=\bm{\beta}^{k}-\eta^{k}\nabla\mathcal{L}(\bm{\beta}^{k}),$
with exponentially decaying step size with base $\eta^{0}=0.01$ scaled by $\sqrt{N/d}$. We consider dimensions $d\in\{10,20,\dots,100\}$, sample sizes $N\in\{100,200,\dots,1000\}$.
Table 2 reports the average squared error $\|\bm{\beta}^{k}-\bm{\beta}^{\star}\|^{2}$ over 1000 replications for both VI and MLE after $k\in\{20,50,100,200\}$ iterations using the softplus link function. The VI estimator consistently attains smaller error than MLE at every iteration count, often by a larger margin especially at the early stage. A more comprehensive set of results with other link functions and a sparse parameter setting can be found in Appendix E, where we observe that VI generally dominates MLE, except for the Gaussian-mixture CDF link in the “easy” regime of large sample size, low dimension, and large iterations.
Table 2: Mean squared error of the VI estimator and MLE with softplus link across iterations $k$. For each $(k,d,N)$ combination, the smaller error between the two estimators is highlighted in bold. The values in the brackets are standard deviations across 1000 independent repetitions.
Link, $k$ $d$ $N=100$ $N=200$ $N=500$ $N=1000$
VI MLE VI MLE VI MLE VI MLE
softplus
$k=20$
10 .627 (.082) .713 (.064) .509 (.070) .616 (.056) .340 (.055) .465 (.047) .212 (.036) .338 (.033)
20 .727 (.065) .791 (.049) .628 (.059) .714 (.046) .469 (.047) .581 (.039) .334 (.037) .461 (.031)
50 .824 (.045) .865 (.034) .753 (.044) .812 (.033) .627 (.036) .713 (.028) .510 (.032) .617 (.025)
100 .881 (.032) .909 (.024) .827 (.032) .868 (.024) .725 (.028) .790 (.022) .627 (.028) .713 (.022)
softplus
$k=50$
10 .467 (.102) .568 (.086) .324 (.081) .439 (.071) .161 (.048) .269 (.048) .076 (.027) .160 (.031)
20 .593 (.090) .674 (.073) .463 (.073) .565 (.061) .281 (.051) .398 (.046) .161 (.033) .270 (.033)
50 .748 (.064) .797 (.050) .637 (.060) .710 (.048) .464 (.049) .566 (.040) .322 (.038) .438 (.033)
100 .829 (.047) .860 (.036) .743 (.046) .794 (.036) .594 (.041) .675 (.033) .461 (.033) .564 (.028)
softplus
$k=100$
10 .401 (.113) .500 (.098) .256 (.080) .363 (.075) .115 (.041) .202 (.044) .049 (.021) .107 (.027)
20 .539 (.099) .621 (.082) .397 (.079) .498 (.068) .215 (.050) .320 (.048) .115 (.030) .202 (.033)
50 .708 (.070) .758 (.056) .585 (.064) .660 (.052) .397 (.049) .498 (.042) .257 (.035) .364 (.033)
100 .815 (.053) .841 (.042) .716 (.052) .764 (.041) .543 (.045) .625 (.037) .397 (.035) .498 (.030)
softplus
$k=200$
10 .384 (.115) .480 (.102) .239 (.080) .341 (.076) .102 (.039) .182 (.044) .045 (.019) .094 (.024)
20 .527 (.105) .607 (.087) .375 (.078) .474 (.068) .198 (.044) .298 (.044) .102 (.028) .182 (.031)
50 .708 (.073) .754 (.058) .571 (.069) .644 (.057) .376 (.049) .476 (.043) .238 (.036) .341 (.034)
100 .815 (.056) .838 (.044) .702 (.055) .750 (.043) .524 (.046) .605 (.038) .376 (.036) .475 (.032)
Figure 3 and Figure 4 illustrate representative convergence behavior for $(d,N)=(20,400)$. Figure 3 shows the sample convergence trajectories of the VI and MLE iterates for all four link functions. The trajectories for the log link coincide exactly, as expected from their equivalence as in (3.1). For non-canonical links, the VI update exhibits noticeably faster convergence in the early iteration phase. In the clipped exponential case, the VI and MLE trajectories coincide during the initial iterations because the iterates have not yet reached the truncation boundary, and the model effectively behaves like the log link in that region. For the Gaussian-mixture CDF link, which induces a highly non-convex regime, VI avoids the oscillations and occasional divergence observed in MLE.
Figure 4 reports the average squared error against the iteration count $k$. The gap between VI and MLE is particularly pronounced for the softplus link, where VI converges significantly faster. For the clipped exponential link, the gap becomes more pronounced in the later iterations as the iterations approach the truncation boundary. For the Gaussian-mixture CDF link, the gap is largest during the early iterations and gradually narrows as both methods stabilize.
Overall, the results demonstrate that the proposed VI framework achieves comparable estimation accuracy to MLE under finite sample sizes, while exhibiting faster convergence and greater numerical stability for non-canonical link functions. With canonical links, both methods coincide, but in more general GLMs the VI formulation effectively mitigates the flat-landscape and conditioning issues that slow gradient-based likelihood estimation.
((a)) log
((b)) softplus
((c)) clipped exponential
((d)) GMM CDF
Figure 3: Convergence trajectories of VI and MLE for Poisson regression with different link functions ($d=20,N=400$).
((a)) softplus
((b)) clipped exponential
((c)) GMM CDF
Figure 4: Average squared error for VI and MLE against iteration budgets for Poisson regression with different link functions ($d=20,N=400$).
7 Discussion
This work reveals the geometry of the variational inequality and offers a new design principle for constructing statistically efficient and computationally stable estimators beyond the classical MLE paradigm. The theoretical insights developed in this work reveal that the convergence and statistical properties of the VI estimator are primarily governed by two quantities: the strong Minty modulus and the Lipschitz constant, which captures local growth around the solution. Together, these constants determine the statistical rate and numerical stability of iterative algorithms. These results suggest a principled strategy for designing link functions that enhance the performance of the VI-based estimator: seek mappings with a larger strong Minty constant and a smaller Lipschitz constant. Such a design improves both convergence speed and asymptotic accuracy. In contrast, MLE-based formulations are more restrictive, since the composite loss $\ell\circ g^{-1}$ must remain consistent with a valid likelihood model. This highlights a key advantage of the VI framework: it permits tailoring the link function directly to the operator geometry without sacrificing interpretability or theoretical soundness. Finally, an important direction for future work is to analyze the behavior of the VI estimator under model misspecification. Understanding the bias-robustness trade-offs induced by non-canonical link functions could provide deeper insight into how operator geometry governs both estimation accuracy and model resilience.
Acknowledgment
This work is partially supported by an NSF DMS-2220495, CNS-2220387, NSF DMS-2134037, and the Coca-Cola Foundation.
Appendix A Examples of Vector Fields for Canonical MLE and VI
This section provides several representative examples of empirical vector fields arising from the MLE and the corresponding VI formulations. For each GLM, we derive the gradient of the empirical NLL function $\mathcal{L}_{N}(\bm{\beta})$ and verify that it coincides with the VI operator $V_{N}(\bm{\beta})$ under the canonical link function.
A.1 Linear regression
For the Gaussian (linear) model with an identity link, the empirical NLL is
$\mathcal{L}_{N}(\bm{\beta})=\frac{1}{N}\sum_{i=1}^{N}\left(g^{-1}(\bm{\beta}^{\top}\tilde{\bm{x}}^{i})-y^{i}\right)^{2},$
whose gradient is
$\nabla\mathcal{L}_{N}(\bm{\beta})=\frac{1}{N}\sum_{i=1}^{N}(g^{-1})^{\prime}(\bm{\beta}^{\top}\tilde{\bm{x}}^{i})\left(g^{-1}(\bm{\beta}^{\top}\tilde{\bm{x}}^{i})-y^{i}\right)\tilde{\bm{x}}^{i}.$
With the identity link $g^{-1}(z)=z$, we have $(g^{-1})^{\prime}(z)=1$, leading to
$\nabla\mathcal{L}_{N}(\bm{\beta})=\frac{1}{N}\sum_{i=1}^{N}\left(g^{-1}(\bm{\beta}^{\top}\tilde{\bm{x}}^{i})-y^{i}\right)\tilde{\bm{x}}^{i}=V_{N}(\bm{\beta}).$
A.2 Exponential regression
For the exponential model, the empirical NLL with $\bm{\beta}^{\top}\tilde{\bm{x}}^{i}>0$ is
$\mathcal{L}_{N}(\bm{\beta})=\frac{1}{N}\sum_{i=1}^{N}\left(\log g^{-1}(\bm{\beta}^{\top}\tilde{\bm{x}}^{i})+y^{i}/g^{-1}(\bm{\beta}^{\top}\tilde{\bm{x}}^{i})\right),$
and the corresponding gradient is
$\nabla\mathcal{L}_{N}(\bm{\beta})=\frac{1}{N}\sum_{i=1}^{N}(g^{-1})^{\prime}(\bm{\beta}^{\top}\tilde{\bm{x}}^{i})\left(\frac{1}{g^{-1}(\bm{\beta}^{\top}\tilde{\bm{x}}^{i})}-\frac{y^{i}}{g^{-1}(\bm{\beta}^{\top}\tilde{\bm{x}}^{i})^{2}}\right)\tilde{\bm{x}}^{i}.$
For the inverse link $g^{-1}(z)=z^{-1}$, we have $(g^{-1})^{\prime}(z)=-z^{-2}$, yielding
$\nabla\mathcal{L}_{N}(\bm{\beta})=-\frac{1}{N}\sum_{i=1}^{N}\left(g^{-1}(\bm{\beta}^{\top}\tilde{\bm{x}}^{i})-y^{i}\right)\tilde{\bm{x}}^{i}=-V_{N}(\bm{\beta}).$
The gradient $\nabla\mathcal{L}_{N}(\bm{\beta})$ differs from $V_{N}(\bm{\beta})$ only by a sign, so minimizing the NLL is equivalent to solving $V_{N}(\bm{\beta})=\mathbf{0}$, yielding the same optimal solution.
A.3 Poisson regression
For the Poisson model, the empirical NLL is
$\mathcal{L}_{N}(\bm{\beta})=\frac{1}{N}\sum_{i=1}^{N}\left(-y^{i}\log g^{-1}(\bm{\beta}^{\top}\tilde{\bm{x}}^{i})+g^{-1}(\bm{\beta}^{\top}\tilde{\bm{x}}^{i})\right).$
The gradient is
$\nabla\mathcal{L}_{N}(\bm{\beta})=\frac{1}{N}\sum_{i=1}^{N}(g^{-1})^{\prime}(\bm{\beta}^{\top}\tilde{\bm{x}}^{i})\left(-\frac{y^{i}}{g^{-1}(\bm{\beta}^{\top}\tilde{\bm{x}}^{i})}+1\right)\tilde{\bm{x}}^{i}.$
With the exponential link $g^{-1}(z)=e^{z}$, we have $(g^{-1})^{\prime}(z)=g^{-1}(z)$, so
$\nabla\mathcal{L}_{N}(\bm{\beta})=\frac{1}{N}\sum_{i=1}^{N}\left(g^{-1}(\bm{\beta}^{\top}\tilde{\bm{x}}^{i})-y^{i}\right)\tilde{\bm{x}}^{i}=V_{N}(\bm{\beta}).$
Across these canonical models, the gradient of the empirical NLL coincides exactly with the VI operator, confirming that the MLE can be viewed as a special case of the VI formulation. This equivalence provides a unified interpretation of estimation problems under different exponential-family link functions.
Appendix B Characterization of Minty Conditions
To start with, we introduce the following definition, which is widely used for solving variational inequalities, e.g., [huang2023beyond].
Definition B.1 (Strong Minty condition).
The vector field $F:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ satisfies the strong Minty condition with constant $\mu\geq 0$ if $\operatorname{Sol}(F)\neq\emptyset$ and
$\langle F(\bm{\beta}),\bm{\beta}-\operatorname{proj}_{\operatorname{Sol}(F)}(\bm{\beta})\rangle\geq\mu\cdot\operatorname{dist}^{2}(\bm{\beta},\operatorname{Sol}(F)),\quad\forall\bm{\beta}\in\mathbb{R}^{d}.$
The strong Minty condition plays a central role in the VI estimator. In this section, we first establish sufficient conditions under which this property holds and illustrate them through representative examples. We then discuss guiding principles for designing link functions that naturally satisfy or approximate the strong Minty condition in practice.
Definition B.2.
Given constants $\rho,\mu_{\mathrm{EB}}>0$, we introduce the following properties:
  • $F$ is said to be $\rho$-weakly monotone if
    $\langle F(\bm{\beta}_{2})-F(\bm{\beta}_{1}),\,\bm{\beta}_{2}-\bm{\beta}_{1}\rangle\;\geq\;-\frac{\rho}{2}\|\bm{\beta}_{2}-\bm{\beta}_{1}\|^{2},\qquad\forall\,\bm{\beta}_{1},\bm{\beta}_{2}\in\mathbb{R}^{d}.$
  • $F$ is said to satisfy the error bound property with constant $\mu_{\mathrm{EB}}>0$ if
    $\|F(\bm{\beta})-F(\operatorname{proj}_{\operatorname{Sol}(F)}(\bm{\beta}))\|\;\geq\;\mu_{\mathrm{EB}}\cdot\mathrm{dist}(\bm{\beta},\,\operatorname{Sol}(F)),\qquad\forall\,\bm{\beta}\in\mathbb{R}^{d}.$
Weak monotonicity is a mild regularity condition that allows certain non-monotone vector fields while still ensuring local stability of the VI solution. In contrast, the error bound condition provides a geometric separation property: it quantifies how far a point $\bm{\beta}$ is from the solution set in terms of the magnitude of the operator, implying that the vector field $F(\bm{\beta})$ is uniformly separated from zero along the direction of its projection onto the solution set. Intuitively, this means that $F$ grows at least linearly with the distance to the solution, forming a local conic region around the zero point. With these two conditions, we have the following implication.
Proposition B.3.
Suppose that the vector field $F:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ satisfies $L$-Lipschitz condition, i.e.,
$\|F(\bm{\beta}_{2})-F(\bm{\beta}_{1})\|\leq L\|\bm{\beta}_{2}-\bm{\beta}_{1}\|.$
Additionally, we assume that $F$ satisfies the $\rho$-weakly monotone condition and the error bound property holds with $\mu_{\rm EB}$. If $\rho L<\mu_{\rm EB}^{2}$, then the strong Minty condition holds with $\mu=\frac{\mu_{\rm EB}^{2}-\rho L}{L-\rho}$.
Proof.
Since $F$ satisfies the $\rho$-weakly monotone and $L$-Lipschitz condition, we know that $F(\bm{\beta})+\rho\bm{\beta}$ is monotone and $L+\rho$ Lipschitz, then by Baillon-Haddad Lemma [baillon1977quelques, Corollaire 10] we know that
$\langle F(\bm{\beta})-F(\operatorname{proj}_{\operatorname{Sol}(F)}(\bm{\beta}))+\rho(\bm{\beta}-\operatorname{proj}_{\operatorname{Sol}(F)}(\bm{\beta})),\bm{\beta}-\operatorname{proj}_{\operatorname{Sol}(F)}(\bm{\beta})\rangle$
$\geq\$$\frac{1}{L+\rho}\|F(\bm{\beta})-F(\operatorname{proj}_{\operatorname{Sol}(F)}(\bm{\beta}))+\rho(\bm{\beta}-\operatorname{proj}_{\operatorname{Sol}(F)}(\bm{\beta}))\|^{2}$
which implies that
$\frac{L-\rho}{L+\rho}\langle F(\bm{\beta})-F(\operatorname{proj}_{\operatorname{Sol}(F)}(\bm{\beta})),\bm{\beta}-\operatorname{proj}_{\operatorname{Sol}(F)}(\bm{\beta})\rangle$
$\geq\$$\frac{1}{L+\rho}\|F(\bm{\beta})-F(\operatorname{proj}_{\operatorname{Sol}(F)}(\bm{\beta}))\|^{2}-\frac{\rho L}{L+\rho}\|\bm{\beta}-\operatorname{proj}_{\operatorname{Sol}(F)}(\bm{\beta})\|^{2}.$
Combining this with the error bound property, we know that
$\langle F(\bm{\beta})-F(\operatorname{proj}_{\operatorname{Sol}(F)}(\bm{\beta})),\bm{\beta}-\operatorname{proj}_{\operatorname{Sol}(F)}(\bm{\beta})\rangle$
[…]
The proof is complete. ∎
It is straightforward to verify that if $F$ is strongly monotone, then it implies the error bound property, which characterizes the local growth behavior around the solution set and implies the strong Minty condition. Moreover, we provide an example showing that $F$ need not be monotone for the strong Minty condition to hold.
Example.
In many practical systems, such as those encountered in signal processing and communications, the ideal input-output response is approximately linear. However, due to hardware imperfections or nonlinear circuit effects, small nonlinear distortions may arise. A simple model capturing such behavior is given by
$g^{-1}(z)=z+2\sin(z)\cos(z).$
The derivative of this function is
$(g^{-1})^{\prime}(z)=1+2(\cos^{2}(z)-\sin^{2}(z))=3\cos^{2}(z)-\sin^{2}(z),$
which is not always positive; hence, $g^{-1}$ is not a monotone function. Nevertheless, let $z^{\star}=0$ denote the solution to $g^{-1}(z)=0$. Then
$\langle g^{-1}(z)-g^{-1}(z^{\star}),\,z-z^{\star}\rangle=g^{-1}(z)z\geq z^{2}+2z\sin(z)\cos(z)=z(z+\sin(2z))\geq\frac{1}{2}z^{2}.$
This inequality confirms that, despite the lack of monotonicity, the mapping $g^{-1}$ satisfies the strong Minty condition with modulus $\mu=1/2$.
Appendix C Proofs for Section 4
Proof of Lemma 4.1
Proof.
Recall the definition that $V_{N}(\bm{\beta})=\frac{1}{N}\sum_{i=1}^{N}(g^{-1}(\beta_{0}+\sum_{j=1}^{d}\beta_{j}x^{i}_{j})-y^{i})\cdot\tilde{\bm{x}}^{i}$. This together with (4.1) implies that
$\langle V_{N}(\bm{\beta}),\bm{\beta}-\operatorname{proj}_{\operatorname{Sol}(V_{N})}(\bm{\beta})\rangle$ $=\left\langle\frac{1}{N}\sum_{i=1}^{N}\left(g^{-1}\left(\beta_{0}+\sum_{j=1}^{d}\beta_{j}x^{i}_{j}\right)-y^{i}\right)\cdot\tilde{\bm{x}}^{i},\bm{\beta}-\operatorname{proj}_{\operatorname{Sol}(V_{N})}(\bm{\beta})\right\rangle$
$\geq\frac{\mu_{g}}{N}\cdot\sum_{i=1}^{N}\left[\left|\tilde{\bm{x}}^{i\top}(\bm{\beta}-\operatorname{proj}_{\operatorname{Sol}(V_{N})}(\bm{\beta}))\right|^{2}\right]$
$=\frac{\mu_{g}}{N}\cdot\left\|\tilde{\bm{X}}_{N}(\bm{\beta}-\operatorname{proj}_{\operatorname{Sol}(V_{N})}(\bm{\beta}))\right\|^{2}$
$\geq\frac{\mu_{g}\sigma_{N}^{2}}{N}\cdot\|\bm{\beta}-\operatorname{proj}_{\operatorname{Sol}(V_{N})}(\bm{\beta})\|^{2}.$
This establishes that the averaged strong Minty condition (4.1) indeed implies the desired inequality for $V_{N}$. It remains to show that (4.1) itself is guaranteed whenever the inverse link function $g^{-1}$ is $\mu_{g}$-strongly monotone. In this case, for any $\bm{\beta}\in\mathbb{R}^{d+1}$ we have
$\sum_{i=1}^{N}\left\langle g^{-1}(\tilde{\bm{x}}^{i\top}\bm{\beta})-y^{i},\tilde{\bm{x}}^{i\top}(\bm{\beta}-\operatorname{proj}_{\operatorname{Sol}(V_{N})}(\bm{\beta}))\right\rangle$
[…]
where the equality is from the definition of $\operatorname{proj}_{\operatorname{Sol}(V_{N})}(\bm{\beta})$, namely that it is a solution of
$V_{N}(\operatorname{proj}_{\operatorname{Sol}(V_{N})}(\bm{\beta}))=\frac{1}{N}\sum_{i=1}^{N}(g^{-1}(\tilde{\bm{x}}^{i\top}\operatorname{proj}_{\operatorname{Sol}(V_{N})}(\bm{\beta}))-y^{i})\cdot\tilde{\bm{x}}^{i}=\mathbf{0}.$
Then (4.1) holds. The proof is complete. ∎
Proof of Lemma 4.2
Proof.
First, from Assumptions 4.1 (ii)-(iii) it follows that $\|\bm{x}\|_{\infty}\leq M$, $\|\bm{x}\|\leq\sqrt{d}M$ and
$\left|g^{-1}\left(\beta_{0}+\sum_{j=1}^{d}\beta_{j}x_{j}\right)-y\right|\leq R.$
Then we have the upper bound (assuming $M\geq 1$) that
$\|V_{(\bm{x},y)}(\bm{\beta})\|_{\infty}$ $\leq\left|g^{-1}\left(\beta_{0}+\sum_{j=1}^{d}\beta_{j}x_{j}\right)-y\right|\cdot\|\tilde{\bm{x}}\|_{\infty}\leq RM,$
$\|V_{(\bm{x},y)}(\bm{\beta})\|$ $\leq\left|g^{-1}\left(\beta_{0}+\sum_{j=1}^{d}\beta_{j}x_{j}\right)-y\right|\cdot\|\tilde{\bm{x}}\|\leq\sqrt{d+1}RM.$
Next, for the Lipschitz constant of $V_{(\bm{x},y)}$, one has that
$\|V_{(\bm{x},y)}(\bm{\beta}^{\prime})-V_{(\bm{x},y)}(\bm{\beta})\|$ $=\left\|\left(g^{-1}\left(\beta_{0}^{\prime}+\sum_{j=1}^{d}\beta^{\prime}_{j}x_{j}\right)-g^{-1}\left(\beta_{0}+\sum_{j=1}^{d}\beta_{j}x_{j}\right)\right)\cdot\tilde{\bm{x}}\right\|$
$\leq L\|\tilde{\bm{x}}\|\cdot\left|\beta_{0}^{\prime}-\beta_{0}+\sum_{j=1}^{d}\beta^{\prime}_{j}x_{j}-\sum_{j=1}^{d}\beta_{j}x_{j}\right|$
$\leq L\|\tilde{\bm{x}}\|^{2}\|\bm{\beta}^{\prime}-\bm{\beta}\|$
$\leq(LdM^{2}+L)\cdot\|\bm{\beta}^{\prime}-\bm{\beta}\|.$
The proof is complete. ∎
Proof of Theorem 4.3
Proof.
First, from the strong Minty condition in Assumption 4.1 (iv), we know that
$\mu\|\bm{\beta}^{\star}-\hat{\bm{\beta}}_{N}\|^{2}\leq\langle V_{N}(\bm{\beta}^{\star}),\bm{\beta}^{\star}-\hat{\bm{\beta}}_{N}\rangle\leq\|V_{N}(\bm{\beta}^{\star})\|\|\bm{\beta}^{\star}-\hat{\bm{\beta}}_{N}\|,$
which implies
$\mu\|\bm{\beta}^{\star}-\hat{\bm{\beta}}_{N}\|\leq\|V_{N}(\bm{\beta}^{\star})\|.$ (C.1)
Recall the definition that for any $\bm{\beta}\in\mathbb{R}^{d+1}$
$V_{N}(\bm{\beta})=\mathbb{E}_{(\bm{x},y)\sim\mathbb{P}_{N}}\left[\left(g^{-1}\left(\beta_{0}+\sum_{j=1}^{d}\beta_{j}x_{j}\right)-y\right)\cdot\tilde{\bm{x}}\right]=\mathbb{E}_{(\bm{x},y)\sim\mathbb{P}_{N}}V_{(\bm{x},y)}(\bm{\beta}).$
Then we know from Lemma 4.2 that
$\|V_{N}(\bm{\beta}^{\star})\|_{\infty}\leq RM.$
As $V(\bm{\beta}^{\star})=\mathbf{0}$, by applying the Hoeffding’s inequality [vershynin2018high, Theorem 2.2.6], it follows that for any $t>0$ and $j\in\{0,\dots,d\}$,
$\mathbb{P}\left(|V_{N}(\bm{\beta}^{\star})_{j}|\geq t\right)$ $=\mathbb{P}\left(|V_{N}(\bm{\beta}^{\star})_{j}-V(\bm{\beta}^{\star})_{j}|\geq t\right)$
$=\mathbb{P}\left(V_{N}(\bm{\beta}^{\star})_{j}-V(\bm{\beta}^{\star})_{j}\geq t\right)+\mathbb{P}\left(-V_{N}(\bm{\beta}^{\star})_{j}+V(\bm{\beta}^{\star})_{j}\geq t\right)$
$\leq 2\exp\left(-\frac{Nt^{2}}{2R^{2}M^{2}}\right).$
Using the union bound (i.e., Boole’s inequality),
$\mathbb{P}\left(\|V_{N}(\bm{\beta}^{\star})\|\geq t\right)$ $\leq\mathbb{P}\left(\sqrt{d+1}\|V_{N}(\bm{\beta}^{\star})\|_{\infty}\geq t\right)$
$\leq\sum_{j=0}^{d}\mathbb{P}\left(|V_{N}(\bm{\beta}^{\star})_{j}|\geq\frac{t}{\sqrt{d+1}}\right)\leq 2(d+1)\exp\left(-\frac{Nt^{2}}{2(d+1)R^{2}M^{2}}\right).$
Equivalently, for any $\epsilon\in(0,1)$, set $t=RM\sqrt{2(d+1)\ln(2(d+1)/\epsilon)/N}$. Then with probability at least $1-\epsilon$,
$\|V_{N}(\bm{\beta}^{\star})\|\leq RM\sqrt{\frac{2(d+1)\ln(2(d+1)/\epsilon)}{N}}.$
Combining this with (C.1), we derive the desired results. ∎
Proof of Lemma 4.4
Proof.
Let $u:=\langle\tilde{\bm{x}},\bm{\beta}^{\star}\rangle$ and $\delta_{t}:=\langle\tilde{\bm{x}},\bm{h}_{t}\rangle$ for some sequence $\bm{h}_{t}\to\bm{h}$ as $t\downarrow 0$. Then
$\frac{V(\bm{\beta}^{\star}+t\bm{h}_{t})-V(\bm{\beta}^{\star})}{t}=\mathbb{E}_{(\bm{x},y)\sim\mathbb{P}}\left[\frac{g^{-1}(u+t\delta_{t})-g^{-1}(u)}{t}\,\tilde{\bm{x}}\right$]. (C.2)
Since $\mathbb{P}(u\in\mathcal{K})=0$ by Assumption 4.2, $g^{-1}$ is differentiable at $u$ almost surely, and therefore
$g^{-1}(u+t\delta_{t})=g^{-1}(u)+(g^{-1})^{\prime}(u)\,t\delta_{t}+o(t|\delta_{t}|)\quad\text{a.s.}$
Hence, pointwise almost surely,
$\frac{g^{-1}(u+t\delta_{t})-g^{-1}(u)}{t}\,\tilde{\bm{x}}\ \rightarrow\ (g^{-1})^{\prime}(u)\,\langle\tilde{\bm{x}},\bm{h}\rangle\,\tilde{\bm{x}}$. (C.3)
Moreover, by the global Lipschitz property of $g^{-1}$ with constant $L$,
$\left\|\frac{g^{-1}(u+t\delta_{t})-g^{-1}(u)}{t}\,\tilde{\bm{x}}\right\|\leq L\,|\delta_{t}|\,\|\tilde{\bm{x}}\|\leq L\,\|\tilde{\bm{x}}\|^{2}\|\bm{h}_{t}\|.$
Under Assumption 4.1, we have $\|\tilde{\bm{x}}\|^{2}\leq 1+dM^{2}$, so the right-hand side is bounded by $L(1+dM^{2})\|\bm{h}_{t}\|$, which is integrable and independent of $t$ since $\|\bm{h}_{t}\|$ is bounded. From (C.2) and (C.3) with the dominated convergence theorem, it follows that
$\frac{V(\bm{\beta}^{\star}+t\bm{h}_{t})-V(\bm{\beta}^{\star})}{t}\ \rightarrow\ \mathbb{E}_{(\bm{x},y)\sim\mathbb{P}}\left[(g^{-1})^{\prime}(\langle\tilde{\bm{x}},\bm{\beta}^{\star}\rangle)\tilde{\bm{x}}\tilde{\bm{x}}^{\top}\right]\bm{h}.$
Thus $V$ is Fréchet differentiable at $\bm{\beta}^{\star}$ with gradient $\nabla V(\bm{\beta}^{\star})$ as claimed. ∎
Proof of Theorem 4.5
Proof.
By Assumption 4.1, $|g^{-1}(\langle\tilde{\bm{x}},\bm{\beta}^{\star}\rangle)-y|\leq R$, $\|\bm{x}\|^{2}\leq dM^{2}$, and $g^{-1}$ is Lipschitz. Hence,
$\mathbb{E}_{(\bm{x},y)\sim\mathbb{P}}\!\left[\|V_{(\bm{x},y)}(\bm{\beta}^{\star})\|^{2}\right]\leq\mathbb{E}_{(\bm{x},y)\sim\mathbb{P}}\!\left[|g^{-1}(\langle\tilde{\bm{x}},\bm{\beta}^{\star}\rangle)-y|^{2}\,\|\tilde{\bm{x}}\|^{2}\right]\leq R^{2}(1+dM^{2})<\infty.$
Therefore, when $(\bm{x}^{i},y^{i})_{i=1}^{N}$ are i.i.d. samples, the random vectors $V_{(\bm{x}^{i},y^{i})}(\bm{\beta}^{\star})$ are i.i.d. with mean zero and covariance matrix $\Gamma$ (finite and positive definite by Assumption (A2)). Applying the multivariate central limit theorem yields
$\sqrt{N}\big(V_{N}(\bm{\beta}^{\star})-V(\bm{\beta}^{\star})\big)\ \xrightarrow{d}\ \mathcal{N}(\mathbf{0},\,\Gamma).$ (C.4)
Adding and subtracting $V(\hat{\bm{\beta}}_{N})$ and $V(\bm{\beta}^{\star})$ on $V_{N}(\hat{\bm{\beta}}_{N})$ yields
$\mathbf{0}=V_{N}(\hat{\bm{\beta}}_{N})$ $=V_{N}(\bm{\beta}^{\star})+\underbrace{V(\hat{\bm{\beta}}_{N})-V(\bm{\beta}^{\star})}_{\text{population increment}}+\underbrace{(V_{N}(\hat{\bm{\beta}}_{N})-V(\hat{\bm{\beta}}_{N}))-(V_{N}(\bm{\beta}^{\star})-V(\bm{\beta}^{\star}))}_{=:R_{N}}.$ (C.5)
By the Fréchet differentiability of $V$ at $\bm{\beta}^{\star}$ (Lemma 4.4), we have the local linearization
$V(\hat{\bm{\beta}}_{N})-V(\bm{\beta}^{\star})=\nabla V(\bm{\beta}^{\star})(\hat{\bm{\beta}}_{N}-\bm{\beta}^{\star})+r_{N},\qquad\|r_{N}\|=o(\|\hat{\bm{\beta}}_{N}-\bm{\beta}^{\star}\|).$ (C.6)
From Lemma 4.2, the sample operator $V_{(\bm{x},y)}$ is $(LdM^{2}+L)$-Lipschitz. Therefore,
$R_{N}=\left[\mathbb{E}_{N}-\mathbb{E}\right]\left(V_{(\cdot)}(\hat{\bm{\beta}}_{N})-V_{(\cdot)}(\bm{\beta}^{\star})\right)$
represents an empirical process evaluated on a Lipschitz-indexed function class. For $0<r\leq 1$, define
$\mathcal{F}_{r}:=\Bigl\{\,f_{\bm{\beta}}(\cdot):=\frac{V_{(\cdot)}(\bm{\beta})-V_{(\cdot)}(\bm{\beta}^{\star})}{\|\bm{\beta}-\bm{\beta}^{\star}\|}\,:\,\|\bm{\beta}-\bm{\beta}^{\star}\|\leq r\Bigr\}.$
Each function $f_{\bm{\beta}}$ is uniformly bounded by the envelope $LdM^{2}+L$, and the class $\mathcal{F}_{r}$ is a finite-dimensional (Euclidean) Lipschitz image of a ball in $\mathbb{R}^{d+1}$. Such finite-dimensional Lipschitz (parametric) classes have a covering numbers that grow polynomially in $1/\epsilon$, a direct consequence of the parameter space being a subset of $\mathbb{R}^{d+1}$ (see [van1996weak, Theorem 2.7.11]). This polynomial entropy bound, together with the existence of a square-integrable envelope, implies that $\mathcal{F}_{r}$ is $\mathbb{P}$-Donsker by [van1996weak, Theorem 2.5.2]. Consequently,
$\sup_{f\in\mathcal{F}_{r}}\big\|\sqrt{N}[\mathbb{E}_{N}-\mathbb{E}]f\big\|=O_{p}(1).$ (C.7)
Since $\|\hat{\bm{\beta}}_{N}-\bm{\beta}^{\star}\|=O_{p}(N^{-1/2})$ by Theorem 4.3, we have
$\mathbb{P}\big(\|\hat{\bm{\beta}}_{N}-\bm{\beta}^{\star}\|\leq r\big)\ \rightarrow\ 1.$
Then on the event $\mathcal{A}_{N}:=\{\|\hat{\bm{\beta}}_{N}-\bm{\beta}^{\star}\|\leq r\}$ it follows
$\sqrt{N}\,\|R_{N}\|=\sqrt{N}\,\left\|[\mathbb{E}_{N}-\mathbb{E}]\left(V_{(\cdot)}(\hat{\bm{\beta}}_{N})-V_{(\cdot)}(\bm{\beta}^{\star})\right)\right\|$ $=\left\|\sqrt{N}[\mathbb{E}_{N}-\mathbb{E}]f_{\hat{\bm{\beta}}_{N}}\right\|\cdot\|\hat{\bm{\beta}}_{N}-\bm{\beta}^{\star}\|$
$\leq\sup_{f\in\mathcal{F}_{r}}\left\|\sqrt{N}[\mathbb{E}_{N}-\mathbb{E}]f\right\|\cdot\|\hat{\bm{\beta}}_{N}-\bm{\beta}^{\star}\|.$
This together with $\|\hat{\bm{\beta}}_{N}-\bm{\beta}^{\star}\|=O_{p}(N^{-1/2})$ and (C.7), we conclude on $\mathcal{A}_{N}$ that
$\sqrt{N}\,\|R_{N}\|=O_{p}(1)\cdot O_{p}(N^{-1/2})=o_{p}(1).$
Since $\mathbb{P}(\mathcal{A}_{N}^{c})\to 0$, the same conclusion holds unconditionally:
$\|R_{N}\|=o_{p}(N^{-1/2}).$
Substitute (C.6) into (C.5):
$\mathbf{0}=V_{N}(\bm{\beta}^{\star})+\nabla V(\bm{\beta}^{\star})(\hat{\bm{\beta}}_{N}-\bm{\beta}^{\star})+r_{N}+R_{N}.$
Rearranging terms gives
$\sqrt{N}(\hat{\bm{\beta}}_{N}-\bm{\beta}^{\star})=-\,\nabla V(\bm{\beta}^{\star})^{-1}\sqrt{N}\big(V_{N}(\bm{\beta}^{\star})-V(\bm{\beta}^{\star})\big)-\nabla V(\bm{\beta}^{\star})^{-1}\sqrt{N}(r_{N}+R_{N}).$
The first term converges in distribution to $\mathcal{N}(0,\nabla V(\bm{\beta}^{\star})^{-1}\Gamma\nabla V(\bm{\beta}^{\star})^{-\top})$ by (C.4). On the other hand, $\sqrt{N}\|r_{N}\|=o_{p}(1)$ and $\sqrt{N}\|R_{N}\|=o_{p}(1)$, the second term vanishes in probability. Applying Slutsky’s theorem gives
$\sqrt{N}(\hat{\bm{\beta}}_{N}-\bm{\beta}^{\star})\ \xrightarrow{d}\ \mathcal{N}\left(\mathbf{0},\nabla V(\bm{\beta}^{\star})^{-1}\Gamma\nabla V(\bm{\beta}^{\star})^{-\top}\right).$
The proof is complete. ∎
Appendix D Proofs for Section 5
Proof of Theorem 5.1
Proof.
From the update rule of the fixed-point iterative method, the iterates satisfy
$\bm{\beta}^{t+1}=\bm{\beta}^{t}-\eta\,V_{N}(\bm{\beta}^{t}),\qquad\hat{\bm{\beta}}_{N}=\hat{\bm{\beta}}_{N}-\eta\,V_{N}(\hat{\bm{\beta}}_{N}),$
where $\hat{\bm{\beta}}_{N}$ is a fixed point of the operator $V_{N}$, i.e., $V_{N}(\hat{\bm{\beta}}_{N})=\mathbf{0}$. It then follows that
$\|\bm{\beta}^{t+1}-\hat{\bm{\beta}}_{N}\|^{2}=\|\bm{\beta}^{t}-\eta\,V_{N}(\bm{\beta}^{t})-\hat{\bm{\beta}}_{N}+\eta\,V_{N}(\hat{\bm{\beta}}_{N})\|^{2}.$
Expanding the square gives
$\|\bm{\beta}^{t+1}-\hat{\bm{\beta}}_{N}\|^{2}=\|\bm{\beta}^{t}-\hat{\bm{\beta}}_{N}\|^{2}-2\eta\big\langle\bm{\beta}^{t}-\hat{\bm{\beta}}_{N},\,V_{N}(\bm{\beta}^{t})-V_{N}(\hat{\bm{\beta}}_{N})\big\rangle+\eta^{2}\|V_{N}(\bm{\beta}^{t})-V_{N}(\hat{\bm{\beta}}_{N})\|^{2}.$
By the strong Minty monotonicity condition in Assumption 4.1 and the Lipschitz continuity in Lemma 4.2, we have
$\left\langle V_{N}(\bm{\beta}^{t})-V_{N}(\hat{\bm{\beta}}_{N}),\bm{\beta}^{t}-\hat{\bm{\beta}}_{N}\right\rangle\geq\mu\|\bm{\beta}^{t}-\hat{\bm{\beta}}_{N}\|^{2},\quad\|V_{N}(\bm{\beta}^{t})-V_{N}(\hat{\bm{\beta}}_{N})\|\leq(LdM^{2}+L)\|\bm{\beta}^{t}-\hat{\bm{\beta}}_{N}\|.$
Substituting these bounds gives
$\|\bm{\beta}^{t+1}-\hat{\bm{\beta}}_{N}\|^{2}\leq\bigl(1-2\eta\mu+\eta^{2}L^{2}(1+dM^{2})^{2}\bigr)\|\bm{\beta}^{t}-\hat{\bm{\beta}}_{N}\|^{2}.$
Choosing $\eta=\mu/(L^{2}(1+dM^{2})^{2})$ yields
$\|\bm{\beta}^{t+1}-\hat{\bm{\beta}}_{N}\|^{2}\leq\left(1-\frac{\mu^{2}}{L^{2}(1+dM^{2})^{2}}\right)\|\bm{\beta}^{t}-\hat{\bm{\beta}}_{N}\|^{2}.$
By applying the above bound recursively, we obtain the claimed linear convergence rate of Algorithm 1. Combining this result with Theorem 4.3 further yields with the probability at least $1-\epsilon$,
$\|\bm{\beta}^{t}-\bm{\beta}^{\star}\|$ $\leq\|\bm{\beta}^{t}-\hat{\bm{\beta}}_{N}\|+\|\hat{\bm{\beta}}_{N}-\bm{\beta}^{\star}\|$
$\leq\left(1-\frac{\mu^{2}}{L^{2}(1+dM^{2})^{2}}\right)^{t/2}\|\bm{\beta}^{0}-\hat{\bm{\beta}}_{N}\|+\frac{RM}{\mu}\sqrt{\frac{2(d+1)\ln(2(d+1)/\epsilon)}{N}}.$
The proof is complete. ∎
Proof of Theorem 5.2
Proof.
From the update rule we know that
$\|\bm{\beta}^{t+1}-\bm{\beta}^{\star}\|^{2}$ $=\bigl\|\bm{\beta}^{t}-\eta^{t}V_{(\bm{x}^{t},y^{t})}(\bm{\beta}^{t})-\bm{\beta}^{\star}\bigr\|^{2}$
$=\|\bm{\beta}^{t}-\bm{\beta}^{\star}\|^{2}-2\eta^{t}\langle V_{(\bm{x}^{t},y^{t})}(\bm{\beta}^{t}),\,\bm{\beta}^{t}-\bm{\beta}^{\star}\rangle+(\eta^{t})^{2}\|V_{(\bm{x}^{t},y^{t})}(\bm{\beta}^{t})\|^{2}.$
Taking expectations and using the law of total expectation yields
$\frac{1}{2}\mathbb{E}\|\bm{\beta}^{t+1}-\bm{\beta}^{\star}\|^{2}$ $\leq\frac{1}{2}\mathbb{E}\|\bm{\beta}^{t}-\bm{\beta}^{\star}\|^{2}-\eta^{t}\,\mathbb{E}\big[\langle V_{(\bm{x}^{t},y^{t})}(\bm{\beta}^{t}),\bm{\beta}^{t}-\bm{\beta}^{\star}\rangle\big]+\frac{1}{2}(\eta^{t})^{2}\,\mathbb{E}\|V_{(\bm{x}^{t},y^{t})}(\bm{\beta}^{t})\|^{2}.$ (D.1)
By the strong Minty condition in Assumption 4.1, we have
$\mathbb{E}\!\left[\langle V_{(\bm{x}^{t},y^{t})}(\bm{\beta}^{t}),\,\bm{\beta}^{t}-\bm{\beta}^{\star}\rangle\right]\geq\mu\,\mathbb{E}\|\bm{\beta}^{t}-\bm{\beta}^{\star}\|^{2}.$
In addition, Lemma 4.2 implies that $\|V_{(\bm{x}^{t},y^{t})}(\bm{\beta}^{t})\|^{2}\leq(d+1)R^{2}M^{2}$. Substituting these bounds into (D.1) yields
$\frac{1}{2}\mathbb{E}\|\bm{\beta}^{t+1}-\bm{\beta}^{\star}\|^{2}\leq\frac{1}{2}(1-2\mu\eta^{t})\mathbb{E}\|\bm{\beta}^{t}-\bm{\beta}^{\star}\|^{2}+\frac{1}{2}(\eta^{t})^{2}(d+1)R^{2}M^{2}.$
Let $\eta^{t}=1/[\mu(t+1)]$ and assume the initialization satisfies
$\frac{1}{2}\mathbb{E}\|\bm{\beta}^{0}-\bm{\beta}^{\star}\|^{2}\leq\frac{c_{0}(d+1)R^{2}M^{2}}{2\mu^{2}}$
for some constant $c_{0}>1$. We will prove by induction that
$\frac{1}{2}\mathbb{E}\|\bm{\beta}^{t}-\bm{\beta}^{\star}\|^{2}\leq\frac{c_{0}(d+1)R^{2}M^{2}}{2\mu^{2}(t+1)},\qquad t=0,1,\ldots.$
The base case $t=0$ holds by assumption. For the induction step, note that $\mu\eta^{t}=1/(t+1)\leq 1/2$. Then,
$\frac{1}{2}\mathbb{E}\|\bm{\beta}^{t+1}-\bm{\beta}^{\star}\|^{2}$ $\leq\frac{1}{2}(1-2\mu\eta^{t})\mathbb{E}\|\bm{\beta}^{t}-\bm{\beta}^{\star}\|^{2}+\frac{1}{2}(\eta^{t})^{2}(d+1)R^{2}M^{2}$
$\leq\frac{c_{0}(d+1)R^{2}M^{2}}{2\mu^{2}t}\left(1-2\mu\eta^{t}\right)+\frac{1}{2}(\eta^{t})^{2}(d+1)R^{2}M^{2}$
$\leq\frac{c_{0}(d+1)R^{2}M^{2}}{2\mu^{2}t}\left(1-\frac{2}{t+1}\right)+\frac{c_{0}(d+1)R^{2}M^{2}}{2\mu^{2}(t+1)^{2}}$
$=\frac{c_{0}(d+1)R^{2}M^{2}}{2\mu^{2}(t+1)}\left(\frac{t-1}{t}+\frac{1}{t+1}\right)\leq\frac{c_{0}(d+1)R^{2}M^{2}}{2\mu^{2}(t+1)}.$
This completes the induction and thus the proof. ∎
Appendix E Additional Experiment Results
Tables 3 and 4 present additional numerical results for the finite-sample performance of VI and MLE under different parameter structures. Table 3 corresponds to the dense parameter setting $\bm{\beta}^{\star}=d^{-1/2}(1,\dots,1)\in\mathbb{R}^{d}$ as in Section 6. The results for the softplus link were already discussed in Table 2, so here we report the outcomes for the log (sanity check), clipped exponential, and Gaussian-mixture CDF links. Table 4 corresponds to the sparse parameter setting, where $\bm{\beta}^{\star}=(2/\sqrt{5},1/\sqrt{5},0,\dots,0)\in\mathbb{R}^{d}$, and summarizes the results for the softplus, clipped exponential, and Gaussian-mixture CDF links (the log link is omitted). Overall, the results under the sparse parameter setting are qualitatively similar to those under the dense parameter setting. In both cases, the VI estimator consistently outperforms MLE across most configurations, with particularly large improvements for the softplus link.
Table 3: Mean squared error of the VI estimator and MLE across link functions and iterations $k$ with dense parameter $\bm{\beta}^{\star}=d^{-1/2}(1,\dots,1)$. For each $(\text{Link},k,d,N)$ combination, the smaller error between the two estimators is highlighted in bold. The values in the brackets are standard deviations across 1000 independent repetitions.
Link, $k$ $d$ $N=100$ $N=200$ $N=500$ $N=1000$
VI MLE VI MLE VI MLE VI MLE
log
$k=20$
10 .241 (.105) .241 (.105) .108 (.057) .108 (.057) .025 (.014) .025 (.014) .007 (.004) .007 (.004)
20 .384 (.106) .384 (.106) .232 (.074) .232 (.074) .079 (.030) .079 (.030) .023 (.009) .023 (.009)
50 .589 (.082) .589 (.082) .432 (.071) .432 (.071) .226 (.046) .226 (.046) .100 (.025) .100 (.025)
100 .719 (.061) .719 (.061) .586 (.062) .586 (.062) .379 (.049) .379 (.049) .221 (.032) .221 (.032)
log
$k=50$
10 .128 (.069) .128 (.069) .048 (.027) .048 (.027) .013 (.007) .013 (.007) .006 (.003) .006 (.003)
20 .258 (.088) .258 (.088) .124 (.049) .124 (.049) .033 (.012) .033 (.012) .013 (.004) .013 (.004)
50 .485 (.080) .485 (.080) .307 (.063) .307 (.063) .120 (.030) .120 (.030) .045 (.011) .045 (.011)
100 .655 (.065) .655 (.065) .486 (.060) .486 (.060) .249 (.039) .249 (.039) .118 (.021) .118 (.021)
log
$k=100$
10 .104 (.059) .104 (.059) .039 (.021) .039 (.021) .012 (.006) .012 (.006) .006 (.003) .006 (.003)
20 .221 (.078) .221 (.078) .101 (.039) .101 (.039) .028 (.010) .028 (.010) .012 (.004) .012 (.004)
50 .452 (.084) .452 (.084) .269 (.059) .269 (.059) .096 (.025) .096 (.025) .037 (.008) .037 (.008)
100 .633 (.064) .633 (.064) .448 (.059) .448 (.059) .214 (.035) .214 (.035) .095 (.017) .095 (.017)
log
$k=200$
10 .096 (.053) .096 (.053) .037 (.020) .037 (.020) .012 (.006) .012 (.006) .006 (.003) .006 (.003)
20 .207 (.075) .207 (.075) .093 (.037) .093 (.037) .027 (.010) .027 (.010) .013 (.004) .013 (.004)
50 .442 (.081) .442 (.081) .259 (.057) .259 (.057) .089 (.022) .089 (.022) .036 (.008) .036 (.008)
100 .626 (.069) .626 (.069) .439 (.059) .439 (.059) .207 (.035) .207 (.035) .090 (.016) .090 (.016)
clipped exp.
$k=20$
10 .593 (.068) .595 (.066) .484 (.054) .488 (.051) .342 (.042) .346 (.036) .241 (.030) .245 (.024)
20 .696 (.052) .697 (.051) .593 (.047) .595 (.045) .449 (.037) .452 (.034) .340 (.028) .344 (.024)
50 .800 (.040) .801 (.039) .722 (.035) .722 (.034) .593 (.030) .595 (.029) .484 (.024) .486 (.023)
100 .865 (.027) .865 (.027) .802 (.026) .803 (.025) .693 (.024) .694 (.023) .594 (.021) .596 (.020)
clipped exp.
$k=50$
10 .444 (.084) .456 (.076) .334 (.066) .344 (.056) .201 (.048) .208 (.036) .125 (.033) .128 (.022)
20 .571 (.074) .580 (.069) .447 (.061) .457 (.055) .297 (.044) .306 (.037) .199 (.032) .206 (.025)
50 .724 (.055) .729 (.051) .608 (.046) .616 (.042) .448 (.038) .458 (.034) .332 (.031) .342 (.026)
100 .814 (.039) .818 (.036) .722 (.038) .727 (.035) .569 (.033) .578 (.030) .450 (.026) .459 (.024)
clipped exp.
$k=100$
10 .388 (.090) .403 (.081) .273 (.068) .285 (.057) .156 (.048) .161 (.034) .089 (.033) .090 (.021)
20 .519 (.080) .532 (.072) .394 (.065) .409 (.057) .239 (.044) .251 (.037) .154 (.033) .160 (.024)
50 .691 (.060) .698 (.055) .566 (.057) .578 (.052) .392 (.041) .406 (.036) .274 (.032) .287 (.026)
100 .803 (.047) .804 (.043) .691 (.043) .698 (.040) .519 (.034) .532 (.031) .390 (.029) .404 (.025)
clipped exp.
$k=200$
10 .377 (.094) .393 (.083) .255 (.070) .270 (.058) .141 (.046) .145 (.033) .080 (.032) .080 (.020)
20 .502 (.084) .516 (.075) .375 (.068) .391 (.058) .228 (.047) .239 (.038) .142 (.034) .147 (.024)
50 .685 (.066) .692 (.060) .547 (.055) .560 (.050) .375 (.042) .391 (.037) .259 (.033) .272 (.027)
100 .800 (.048) .799 (.043) .680 (.046) .688 (.041) .507 (.038) .521 (.034) .376 (.030) .391 (.026)
GMM CDF
$k=20$
10 .167 (.061) .459 (.094) .076 (.030) .271 (.069) .023 (.010) .074 (.026) .010 (.004) .017 (.007)
20 .302 (.070) .610 (.070) .166 (.043) .456 (.068) .057 (.017) .209 (.043) .022 (.007) .073 (.017)
50 .513 (.067) .754 (.045) .348 (.050) .648 (.044) .164 (.026) .452 (.043) .076 (.013) .267 (.033)
100 .663 (.053) .827 (.031) .511 (.048) .753 (.032) .299 (.031) .609 (.032) .165 (.019) .451 (.029)
GMM CDF
$k=50$
10 .103 (.045) .200 (.079) .049 (.023) .079 (.036) .019 (.009) .017 (.008) .010 (.004) .007 (.003)
20 .210 (.066) .371 (.086) .105 (.035) .201 (.053) .039 (.013) .055 (.018) .019 (.006) .017 (.005)
50 .435 (.074) .588 (.068) .260 (.051) .426 (.059) .104 (.022) .199 (.035) .049 (.010) .079 (.015)
100 .624 (.063) .715 (.048) .435 (.055) .586 (.050) .213 (.028) .370 (.038) .105 (.015) .199 (.025)
GMM CDF
$k=100$
10 .098 (.046) .145 (.066) .048 (.023) .052 (.025) .019 (.009) .014 (.006) .010 (.005) .007 (.003)
20 .200 (.065) .297 (.082) .100 (.033) .143 (.045) .039 (.013) .037 (.013) .019 (.006) .014 (.005)
50 .424 (.079) .525 (.072) .240 (.049) .344 (.058) .099 (.021) .141 (.028) .048 (.010) .051 (.011)
100 .625 (.070) .671 (.056) .419 (.056) .520 (.051) .197 (.027) .292 (.036) .098 (.015) .140 (.020)
GMM CDF
$k=200$
10 .099 (.050) .131 (.067) .049 (.023) .047 (.023) .020 (.010) .014 (.007) .010 (.004) .007 (.003)
20 .196 (.066) .273 (.079) .097 (.032) .127 (.042) .038 (.013) .034 (.012) .019 (.006) .014 (.004)
50 .421 (.083) .507 (.077) .241 (.049) .328 (.057) .098 (.020) .125 (.027) .049 (.010) .047 (.010)
100 .629 (.075) .660 (.057) .421 (.056) .503 (.052) .195 (.029) .271 (.035) .097 (.015) .127 (.020)
Table 4: Mean squared error of the VI estimator and MLE across link functions and iterations $k$ with sparse parameter $\bm{\beta}^{\star}=(2/\sqrt{5},1/\sqrt{5},0,\dots,0)$.
Link, $k$ $d$ $N=100$ $N=200$ $N=500$ $N=1000$
VI MLE VI MLE VI MLE VI MLE
softplus
$k=20$
10 .632 (.078) .717 (.061) .513 (.069) .619 (.055) .339 (.054) .464 (.046) .216 (.037) .342 (.034)
20 .731 (.064) .795 (.049) .625 (.058) .711 (.046) .463 (.047) .577 (.038) .334 (.036) .461 (.030)
50 .832 (.043) .872 (.032) .757 (.042) .815 (.032) .631 (.037) .716 (.028) .508 (.030) .615 (.024)
100 .882 (.031) .909 (.023) .823 (.029) .865 (.022) .726 (.028) .791 (.021) .626 (.027) .712 (.021)
softplus
$k=50$
10 .469 (.106) .570 (.088) .341 (.080) .455 (.069) .159 (.048) .268 (.048) .078 (.028) .161 (.032)
20 .585 (.082) .668 (.066) .459 (.078) .562 (.066) .276 (.053) .393 (.048) .164 (.033) .273 (.033)
50 .748 (.058) .798 (.046) .637 (.062) .710 (.050) .460 (.052) .563 (.043) .324 (.034) .439 (.030)
100 .827 (.046) .858 (.036) .741 (.047) .792 (.037) .596 (.039) .677 (.032) .465 (.032) .567 (.027)
softplus
$k=100$
10 .400 (.117) .498 (.102) .258 (.080) .365 (.075) .115 (.042) .201 (.046) .048 (.021) .105 (.027)
20 .535 (.103) .617 (.085) .401 (.083) .501 (.070) .215 (.051) .321 (.049) .116 (.035) .204 (.038)
50 .717 (.064) .766 (.050) .584 (.063) .659 (.052) .397 (.048) .498 (.041) .258 (.037) .365 (.034)
100 .818 (.053) .843 (.042) .706 (.053) .756 (.043) .539 (.042) .621 (.035) .394 (.031) .495 (.027)
softplus
$k=200$
10 .378 (.111) .476 (.099) .232 (.079) .335 (.077) .100 (.037) .179 (.043) .045 (.019) .093 (.025)
20 .527 (.106) .607 (.089) .382 (.080) .479 (.071) .200 (.051) .300 (.050) .102 (.029) .182 (.032)
50 .707 (.075) .754 (.061) .574 (.069) .647 (.057) .375 (.049) .475 (.043) .241 (.034) .344 (.032)
100 .817 (.057) .839 (.045) .701 (.058) .750 (.046) .521 (.045) .603 (.038) .376 (.036) .475 (.032)
clipped exp.
$k=20$
10 .596 (.064) .598 (.062) .487 (.057) .491 (.054) .339 (.039) .343 (.035) .241 (.030) .244 (.024)
20 .694 (.053) .695 (.051) .595 (.042) .597 (.040) .444 (.038) .448 (.035) .341 (.027) .345 (.024)
50 .800 (.037) .801 (.036) .721 (.036) .722 (.035) .590 (.028) .592 (.027) .485 (.025) .488 (.023)
100 .867 (.024) .868 (.024) .799 (.025) .799 (.025) .692 (.025) .693 (.024) .595 (.019) .597 (.019)
clipped exp.
$k=50$
10 .447 (.085) .457 (.077) .325 (.061) .337 (.051) .203 (.045) .210 (.034) .126 (.036) .130 (.023)
20 .572 (.067) .581 (.061) .445 (.063) .459 (.056) .298 (.045) .308 (.037) .199 (.031) .206 (.024)
50 .720 (.058) .726 (.054) .611 (.048) .619 (.045) .445 (.041) .455 (.036) .332 (.027) .342 (.023)
100 .813 (.040) .816 (.037) .724 (.041) .729 (.038) .566 (.032) .575 (.029) .446 (.027) .457 (.024)
clipped exp.
$k=100$
10 .390 (.093) .406 (.083) .275 (.068) .287 (.055) .156 (.045) .164 (.033) .091 (.036) .091 (.025)
20 .520 (.082) .533 (.074) .391 (.062) .407 (.056) .241 (.049) .251 (.038) .154 (.034) .159 (.025)
50 .684 (.054) .692 (.051) .567 (.052) .578 (.048) .389 (.043) .404 (.037) .278 (.032) .290 (.027)
100 .799 (.046) .801 (.041) .688 (.045) .696 (.041) .521 (.036) .533 (.032) .393 (.030) .407 (.026)
clipped exp.
$k=200$
10 .384 (.085) .396 (.076) .259 (.072) .273 (.061) .142 (.046) .148 (.035) .078 (.030) .080 (.019)
20 .506 (.086) .520 (.078) .368 (.065) .386 (.057) .230 (.048) .239 (.038) .142 (.034) .149 (.024)
50 .676 (.063) .683 (.059) .550 (.054) .562 (.049) .373 (.040) .389 (.034) .265 (.032) .276 (.026)
100 .798 (.054) .798 (.048) .682 (.047) .690 (.043) .504 (.037) .518 (.033) .374 (.031) .390 (.027)
GMM CDF
$k=20$
10 .171 (.066) .467 (.097) .075 (.027) .273 (.068) .022 (.010) .073 (.023) .009 (.004) .017 (.007)
20 .304 (.072) .608 (.070) .164 (.041) .456 (.065) .057 (.017) .208 (.038) .022 (.007) .074 (.016)
50 .512 (.066) .756 (.043) .348 (.052) .647 (.048) .166 (.026) .455 (.040) .075 (.014) .266 (.031)
100 .666 (.051) .828 (.030) .515 (.047) .755 (.034) .302 (.031) .611 (.031) .165 (.019) .453 (.031)
GMM CDF
$k=50$
10 .107 (.054) .201 (.086) .050 (.023) .084 (.037) .019 (.009) .017 (.008) .009 (.004) .007 (.003)
20 .221 (.072) .384 (.093) .105 (.035) .199 (.053) .039 (.013) .054 (.018) .019 (.006) .017 (.005)
50 .421 (.079) .577 (.071) .259 (.050) .429 (.058) .104 (.019) .199 (.033) .048 (.009) .076 (.014)
100 .627 (.070) .719 (.053) .433 (.057) .585 (.050) .212 (.027) .370 (.035) .104 (.015) .199 (.023)
GMM CDF
$k=100$
10 .095 (.046) .134 (.063) .051 (.026) .055 (.027) .019 (.008) .014 (.006) .010 (.005) .006 (.003)
20 .201 (.063) .301 (.087) .098 (.033) .143 (.050) .038 (.012) .037 (.012) .019 (.006) .014 (.004)
50 .421 (.081) .522 (.071) .244 (.049) .347 (.063) .098 (.020) .141 (.029) .049 (.011) .052 (.012)
100 .623 (.068) .670 (.054) .425 (.052) .522 (.049) .199 (.032) .288 (.037) .100 (.015) .142 (.021)
GMM CDF
$k=200$
10 .096 (.047) .128 (.064) .051 (.024) .048 (.023) .021 (.010) .014 (.007) .009 (.004) .007 (.003)
20 .192 (.060) .277 (.082) .095 (.033) .130 (.041) .039 (.013) .033 (.011) .020 (.006) .014 (.005)
50 .415 (.077) .499 (.071) .245 (.050) .331 (.063) .099 (.020) .131 (.026) .048 (.010) .046 (.010)
100 .629 (.076) .659 (.059) .426 (.061) .506 (.056) .196 (.031) .273 (.037) .098 (.015) .125 (.019)